### v2

# A Power Method for Computing Singular Value Decomposition

Abstract
The singular value decomposition (SVD) allows to put a matrix as a product of three matrices: a matrix with the left singular vectors, a matrix with the positive-valued singular values and a matrix with the right singular vectors. There are two main approaches allowing to get the SVD result: the classical method and the randomized method. The analysis of the classical approach leads to accurate singular values. The randomized approach is especially used for high dimensional matrix and is based on the approximation accuracy without computing necessary all singular values. In this paper, the SVD computation is formalized as an optimization problem and a use of the gradient search algorithm. That results in a power method allowing to get all or the first largest singular values and their associated right vectors. In this iterative search, the accuracy on the singular values and the associated vector matrix depends on the user settings. Two applications of the SVD are the principal component analysis and the autoencoder used in the neural network models.
Keywords:
eigendecomposition, power method, singular values, principal components, autoencoder, neural network.
1 Introduction
Nowadays, we are dealing with massive data in our current activities: words count in text documents, image, gene expression from sequencing analyses, etc. Such data are typically organized in a matrix of size $m\times n$ where $m$ and $n$ are respectively the number of rows and columns of this matrix. The values $m$ and $n$ also define the matrix: $\min(m,n)$. We assume that $m\geq n$, this condition can be met by using the matrix transpose. One characteristic of the matrix associated with the words count in a text document, image, single-cell RNA sequencing or similar data, is the number of zero entries which can be high, up to $50\%$ or more. The singular value decomposition (SVD) allows to put a matrix into a product of three matrices. The properties of these factor matrices allow to write the matrix as a sum of rank one matrices, [9]. The contribution of each term in this sum is quantified by a nonnegative value called singular value. The classical SVD computation methods proceed in two steps: bidiagonalization or rank-revealing decomposition of the matrix, then, search for the factor matrices, [13, 14, 6, 8]. Usually, all singular values and associated vectors are computed using a classical method. In some data analysis methods, the smallest singular value and corresponding vectors may be important, [8]. For the various high dimensional data encountered todays, compact form of the data is often required for storing purpose or for revealing some links between data objects. That is why randomized methods for computing SVD come in play. Indeed, the first important singular values are necessary to get some useful information from the data. These methods also proceed into two steps: search for projection feature space, then obtention of the SVD factor matrices, [17, 16].
In this paper, a method for computing SVD is proposed. This algorithm is especially suitable for searching for few singular values of a high dimensional data but it also applied to small size matrices. The proposed algorithm is inspired from the probabilistic random search approaches by [17, 16]. The difference comes from our formulation of the SVD computing as an objective function optimization problem and the use of a power method to get the projection matrix.
1.1 Classical SVD computation
The SVD allows to decompose any matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$ as a product of three other matrices, [9, 28, 15, 18]:
$\mathbf{X}=\mathbf{UD}\mathbf{W}^{\mathsf{T}}$ (1)
where $\mathbf{U}$ and $\mathbf{W}$ are square column orthogonal matrices of order $m$ and $n$, respectively, $\mathbf{D}$ is a matrix of size $m\times n$. Only the diagonal entries of the upper square part of the matrix $\mathbf{D}$ are nonzero and are the positive-valued singular values: $d_{11}\geq d_{22}\geq\ldots\geq d_{nn}$. In the thin SVD, the matrix $\mathbf{U}$ is of size $m\times n$ and $\mathbf{D}$ is a square diagonal matrix of order $n$. Let us observe that $-\mathbf{U}$ and $-\mathbf{W}$ are jointly also solution of the singular value decomposition in relation (1). The column orthogonal property of the matrices $\mathbf{U}$ and $\mathbf{W}$ allows to write: $\mathbf{U}^{\mathsf{T}}\mathbf{U}=\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{n}$, where $\mathbf{I}_{n}$ is the identity matrix of order $n$. Using a sequence of Householder transformations, a bidiagonal form is obtained: $\mathbf{X}=\mathbf{PGDH}^{\mathsf{T}}\mathbf{Q}^{\mathsf{T}}=\mathbf{PBQ}^{\mathsf{T}}$, where the upper square part of the matrix $\mathbf{B}$ is triangular with the same singular values as the matrix $\mathbf{X}$, $\mathbf{U}=\mathbf{PG}$ and $\mathbf{W}=\mathbf{QH}$. The singular values are easily computed from the matrix $\mathbf{B}$ leading to the matrices $\mathbf{U}$ and $\mathbf{W}$. The LDU (lower triangular, diagonal and upper triangular) decomposition is used by [8] to obtain the SVD. The diagonal entries of the triangular matrices $\mathbf{L}$ and $\mathbf{U}$ are all $1$. The diagonalization of the matrix $\mathbf{X}$ before obtaining the SVD leads to relatively accurate singular values. All singular values and their associated vectors are recovered using the classical SVD method.
Using the column orthogonal property of the matrices $\mathbf{U}$ and $\mathbf{W}$, the decomposition in relation (1) can be written as:
$\mathbf{X}=\sum_{j=1}^{n}d_{jj}\mathbf{u}_{j}\mathbf{w}_{j}^{\mathsf{T}}$ (2)
where $d_{jj}$ are the singular values: $d_{11}\geq d_{22}\geq\ldots\geq d_{nn}$, the vectors $\mathbf{u}_{j}$ and $\mathbf{w}_{j}$ are the $j$-th columns of the matrices $\mathbf{U}$ and $\mathbf{W}$, respectively, $\mathbf{u}_{k}^{\mathsf{T}}\mathbf{u}_{j}=1$, if $k=j$, 00 otherwise, idem for vectors $\mathbf{w}_{j}$. Again, for column $j$, the couples of vectors ($\mathbf{u}_{j},\mathbf{w}_{j}$) and ($-\mathbf{u}_{j},-\mathbf{w}_{j}$) have the same contribution. If there are $n$ nonzero scalars $d_{jj}$, then the matrix $\mathbf{X}$ is of full rank. Otherwise, the number of nonzero scalars $d_{jj}$ is equal to the actual rank of the matrix $\mathbf{X}$. The sum relation (2) is often truncated to $r$ terms ($r<n$) leading to an approximate result of rank lower than data matrix. The percentage of the information captured by the $r$ largest singular values is quantified as:
$reconst.(\%)=100\left(\sum_{j=1}^{r}d_{jj}\left/\sum_{j=1}^{n}d_{jj}\right.\right)$ (3)
where the parameter $r$ is chosen to have the desired percentage in data reconstruction applications.
The Frobenius norm of the matrix $\mathbf{X}$ is:
$\|\mathbf{X}\|_{\mathsf{F}}^{2}=\sum_{i=1}^{m}\sum_{j=1}^{n}x_{ij}^{2}=\operatorname{tr}\left[\mathbf{X}^{\mathsf{T}}\mathbf{X}\right]=\operatorname{tr}\left[\mathbf{XX}^{\mathsf{T}}\right]=\sum_{j=1}^{n}d_{jj}^{2}$ (4)
where $\operatorname{tr}[.]$ is the matrix trace operator. Let $\mathbf{X}_{r}$ be the matrix obtained using the first $r$ terms in the sum of relation (2). The Frobenius norm allows to quantify the error of the $r$ rank approximation for $\mathbf{X}$, [15, 18]:
$\|\mathbf{X}-\mathbf{X}_{r}\|_{\mathsf{F}}^{2}=\sum_{j=r+1}^{n}d_{jj}^{2}$ (5)
The SVD also allows to solve the linear system of equations which occurs in many engineering problems: $\mathbf{Ax}=\mathbf{b}$, where $\mathbf{A}$ and $\mathbf{b}$ are known matrix and vector of size $m\times n$ and $n$, respectively. The entries of the unknown vector $\mathbf{x}$ should be found. For a square matrix, $\mathbf{A}$, the Lanczos algorithm, [23], allows to obtain the solution through matrix eigenvalues, see below. For a non-square matrix $\mathbf{A}$, the normal equation, $\mathbf{A}^{\mathsf{T}}\mathbf{Ax}=\mathbf{A}^{\mathsf{T}}\mathbf{b}$ which involves a symmetric can be used to search for the vector $\mathbf{x}$. For a sparse matrix with large $m$ and $n$ values, the Lanczos algorithm has been extended by Larsen, [25], to bidiagonalization of the non-square matrix leading to perform its SVD. For high dimensional data, when the purpose is to get an approximation of the data matrix, it may not be necessary to get all singular values. The Lanczos algorithm extension proposed in [1] allows to compute the few first largest or smallest singular values and their associated vectors.
1.2 Probabilistic SVD computation
Performing complete SVD can be computationally costly, especially for high dimensional data. When the first important singular values are required, the randomized methods may be used. The computation of the SVD for massive data has been considered in [27, 17, 16, 38], where, only the needed $r$ singular values and their associated vectors are computed using a probabilistic (randomized) approach. The search consists of finding column orthonormal matrix $\mathbf{Q}$ of size $m\times r$ such that:
$\|\mathbf{X}-\mathbf{QQ}^{\mathsf{T}}\mathbf{X}\|\leq\epsilon$ (6)
[…]
$(\mathbf{XX}^{\mathsf{T}})^{q}\mathbf{X}=\mathbf{U}\mathbf{D}^{2q+1}\mathbf{W}^{\mathsf{T}}$ (7)
This relation shows a ($12q+$ 1)-power of the singular values for this matrix $\mathbf{Y}$. The rapid singular values decay leads to a fast convergence algorithm, [17, 16]. The paper [17] gives historical overview of the randomized computing and its applications in linear algebra problems. Using the probabilistic SVD method, the rank $r$ approximation error is slightly greater than the bound given in relation (5).
1.3 Eigendecomposition
The eigendecomposition can be performed for any square matrix. Here, only the symmetric matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ case is considered. The eigendecomposition of this matrix is:
$\mathbf{X}^{\mathsf{T}}\mathbf{X}=\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{\mathsf{T}}$ (8)
where $\boldsymbol{\Lambda}$ is a diagonal matrix which entries are nonnegative-valued scalars or eigenvalues: $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}\geq 0$, $\mathbf{V}$ is an orthogonal matrix, $\mathbf{V}^{\mathsf{T}}\mathbf{V}=\mathbf{I}_{n}$, which columns are the eigenvectors associated with the eigenvalues. Like in relation (2), the eigendecomposition can be expressed as the sum of rank one matrices or spectral decomposition:
$\mathbf{X}^{\mathsf{T}}\mathbf{X}=\sum_{j=1}^{n}\lambda_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{\mathsf{T}}$ (9)
where $\mathbf{v}_{j}$ is the $j$-th column vector of the matrix $\mathbf{V}$. The above sum can also be truncated to get an lower rank $r<n$ approximation for the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$. The column orthogonal property of the matrix $\mathbf{V}$ allows to have:
$(\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}=\mathbf{V}\boldsymbol{\Lambda}^{q}\mathbf{V}^{\mathsf{T}}$ (10)
where $q$ is an integer, $q\geq 1$. That means, all power $q$ of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ have the same eigenvectors, only the eigenvalues differ by the power $q$ and we also have: $\lambda_{1}^{q}\geq\lambda_{2}^{q}\geq\ldots\geq\lambda_{n}^{q}\geq 0$. At least, the first eigenvalue is nonzero. Therefore, we have:
$(a)$ ::: $1\geq(\lambda_{2}/\lambda_{1})\geq(\lambda_{3}/\lambda_{1})\geq\ldots\geq(\lambda_{n}/\lambda_{1})$ (11)
$(b)$ ::: $1\geq(\lambda_{2}/\lambda_{1})^{q}\geq(\lambda_{3}/\lambda_{1})^{q}\geq\ldots\geq(\lambda_{n}/\lambda_{1})^{q}$ (12)
$(c)$ ::: $(\lambda_{j}/\lambda_{1})^{q}\leq(\lambda_{j}/\lambda_{1})\mbox{ ; }j=1,2,\ldots,n$ (13)
In the principal component analysis, see below, the columns of the matrix $\mathbf{V}$ correspond to projection subspaces. One can observe that $-\mathbf{V}$ is also solution of the eigendecomposition in relation (8). This observation allows to fulfil the Perron-Frobenius theorem, [31, 12], which stipulates that the entries of the main eigenvector of a nonnegative-valued matrix are all positive or zero. Nonnegative-valued matrices occur for many real-life data, e.g. genes expression and images. The main eigenvector is used in ranking applications, [5, 24]. To compute the eigenvalue with maximum modulus and the associated eigenvector, the iterative power method, [15], is very often used, see algorithm 2 in the Annexes.
The convergence rate of the power method depends on the modulus of the second eigenvalue compared to the first, [15, 7]. At the convergence, the eigenvector is used to get the associated eigenvalue: $\mathbf{X}^{\mathsf{T}}\mathbf{Xv}_{j}=\lambda_{j}\mathbf{v}_{j}$. Thank to relation (9), the power method can be used to obtain all eigenvectors and their associated eigenvalues. Indeed, let us define $\mathbf{S}_{1}=\mathbf{X}^{\mathsf{T}}\mathbf{X}$. When the first eigenvector is obtained, a new matrix can be formed as $\mathbf{S}_{2}=\mathbf{S}_{1}-\lambda_{1}\mathbf{v}_{1}\mathbf{v}_{1}^{\mathsf{T}}$. The matrix $\mathbf{S}_{2}$ is used to get the second eigenvector and associated eigenvalue. This process is repeated to get all the eigenvector and eigenvalue pairs. That is the Jordan SVD algorithm, [37]. For a rank deficient matrix $\mathbf{X}$, the number of eigenvalues is less than $n$. The power method proposed below allows to compute all required eigenvalues together.
2 Computing singular values using a power method
Given a matrix $\mathbf{X}$ of size $m\times n$, $m\geq n$, we first show the link between it’s singular value decomposition factors and the eigendecomposition factors of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$. Using the expression (1) of $\mathbf{X}$ in the relation (8) allows to write:
$\mathbf{X}^{\mathsf{T}}\mathbf{X}=(\mathbf{UDW}^{\mathsf{T}})^{\mathsf{T}}\mathbf{UDW}^{\mathsf{T}}=\mathbf{WD}^{2}\mathbf{W}^{\mathsf{T}}$
Comparing the above expression to relation (8) allows to express the SVD factors using the eigendecomposition factors:
[…]
The entries of the diagonal matrix $\boldsymbol{\Lambda}$ are the diagonal entries of the matrix $(\mathbf{XW})^{\mathsf{T}}\mathbf{XW}$. Relation (14) shows a one-to-one relationship between the SVD factors of the matrix $\mathbf{X}$ and the eigendecomposition factors of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$. From the above relation, the SVD of the matrix $\mathbf{X}$ can be written as:
$\mathbf{X}=\mathbf{XWW}^{\mathsf{T}}$ (15)
When $\mathbf{W}\triangleq\mathbf{Q}$, the above relation is equivalent to that is used in relation (6) and is obtained when performing the eigendecomposition of the matrix $\mathbf{XX}^{\mathsf{T}}$. Since we assumed that $n\leq m$, working with the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ necessitates less computational load compared to the matrix $\mathbf{XX}^{\mathsf{T}}$. To obtain the column orthogonal matrix $\mathbf{W}$ given $\mathbf{X}$, a norm of the difference between the left-hand and right-hand side terms in the above relation can be minimized. In the probabilistic SVD approach, a column orthonormal matrix of size $r+\ell$, $\ell>r$, is used and the QR decomposition is performed followed by matrices multiplication to get the SVD result. Here, the entries of a random column orthogonal matrix of size $n\times r$ are iteratively modified to fit the factor $\mathbf{W}$ for the SVD result of the matrix $\mathbf{X}$. The algorithm used is similar to the power method but all eigenvectors are simultaneously recovered at a time. A new parameter is introduced to increase the convergence rate, compared to the basic power method.
2.1 Search for the projection matrix using the gradient method
To obtain the matrix $\mathbf{W}$ in relation (15), the Frobenius norm of the difference between $\mathbf{X}$ and $\mathbf{XWW}^{\mathsf{T}}$ is minimized, subject to the column orthogonality constraint for the matrix $\mathbf{W}$. The following constrained optimization problem is solved:
$\min\mathsf{J}(\mathbf{W})\mbox{, }\mathsf{J}(\mathbf{W})=\frac{1}{2}\left\|\mathbf{X}-\mathbf{XWW}^{\mathsf{T}}\right\|_{\mathsf{F}}^{2}\mbox{ subject to: }\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}$ (16)
where .$\|.\|_{\mathsf{F}}^{2}$ is the square of the Frobenius norm. The Lagrangian of this optimization problem is:
$\mathcal{L}(\mathbf{W},\Upsilon)=\frac{1}{2}\left\|\mathbf{X}-\mathbf{XWW}^{\mathsf{T}}\right\|_{\mathsf{F}}^{2}-\left\|\Upsilon\left(\mathbf{W}^{\mathsf{T}}\mathbf{W}-\mathbf{I}\right)\right\|_{\mathsf{F}}^{2}$ (17)
where the entries of the matrix $\Upsilon$ are the Lagrange multipliers associated with the equality constraints. To minimize $\mathsf{J}(\mathbf{W})$, the gradient descent method can be used with the column orthogonal constraint for the matrix $\mathbf{W}$. For the gradient search optimization method, the solution at iteration $t$ is obtained after updating the previous one: $\mathbf{W}^{(t)}=\mathbf{W}^{(t-1)}-\eta\nabla\mathcal{L}(\mathbf{W},\Upsilon)$, where $\eta$ is a positive step parameter and $\nabla\mathcal{L}(\mathbf{W},\Upsilon)$ is the gradient of the Lagrangian. To obtain the matrix $\mathbf{W}$ leading to minimum of the Lagrangian, the derivative of $\mathcal{L}(\mathbf{W},\Upsilon)$ with respect to $\mathbf{W}$ and $\Upsilon$ should be computed.
Relation (4) allows to write:
$\left\|\mathbf{X}-\mathbf{XWW}^{\mathsf{T}}\right\|_{\mathsf{F}}^{2}$ $=$ $\operatorname{tr}\left[(\mathbf{X}-\mathbf{XWW}^{\mathsf{T}})(\mathbf{X}-\mathbf{XWW}^{\mathsf{T}})^{\mathsf{T}}\right]$
$=$ $\operatorname{tr}\left[\mathbf{X}\mathbf{X}^{\mathsf{T}}-2\mathbf{XWW}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}+\mathbf{XWW}^{\mathsf{T}}\mathbf{WW}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}\right]$
For any column orthogonal matrix $\mathbf{W}$ of size $n\times r$, $\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{r}$, we then have:
$\left\|\mathbf{X}-\mathbf{XWW}^{\mathsf{T}}\right\|_{\mathsf{F}}^{2}$ $=$ $\operatorname{tr}\left[\mathbf{X}\mathbf{X}^{\mathsf{T}}-\mathbf{XWW}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}\right]$ (18)
$\left\|\mathbf{W}^{\mathsf{T}}\mathbf{W}-\mathbf{I}\right\|_{\mathsf{F}}^{2}$ $=$ 00 0 (19)
For an $n$-order orthogonal matrix $\mathbf{W}$, we also have $\mathbf{W}\mathbf{W}^{\mathsf{T}}=\mathbf{I}_{n}$. However, all $n$-order orthogonal matrices do not allow to verify relation (8) for a given matrix $\mathbf{X}$. In addition, we would like to be able to compute a column orthogonal matrix $\mathbf{W}$ of size $n\times r$, $r<n$. In this case, we still have $\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{r}$ but $\mathbf{W}\mathbf{W}^{\mathsf{T}}\neq\mathbf{I}_{n}$. The matrix $\mathbf{XX}^{\mathsf{T}}-\mathbf{XWW}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}$ is symmetric and has nonnegative-valued eigenvalues. The trace of this matrix is obtained as the sum of its eigenvalues. For an $n$-order column orthogonal matrix $\mathbf{W}$, relation (18) shows that the global minimum of the objective function $\mathsf{J}(\mathbf{W})$ is zero. For an $r$-column orthogonal matrix $\mathbf{W}$, $r<n$, we have:
$tr\left[\mathbf{XX}^{\mathsf{T}}-\mathbf{XWW}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}\right]=tr\left[\mathbf{XX}^{\mathsf{T}}\right]-tr\left[\mathbf{XW}(\mathbf{XW})^{\mathsf{T}}\right]=\sum_{j=r+1}^{n}d_{jj}$
where $\mathbf{XW}(\mathbf{XW})^{\mathsf{T}}$ is the diagonal matrix which entries are the $r$ first singular values, see the text under relation (14).
The relation (118) in Petersen & Petersen [32, page 13], can be written as:
$\frac{\partial}{\partial\mathbf{W}}\operatorname{tr}\left[\mathbf{AWBW}^{\mathsf{T}}\mathbf{C}\right]=\mathbf{A}^{\mathsf{T}}\mathbf{C}^{\mathsf{T}}\mathbf{WB}^{\mathsf{T}}+\mathbf{CAWB}$
Using this relation with $\mathbf{A}=\mathbf{X}$, $\mathbf{B}=\mathbf{I}$ and $\mathbf{C}=\mathbf{X}^{\mathsf{T}}$, the gradient of the objective function $\mathsf{J}(\mathbf{W})$ is:
$\frac{\partial\mathsf{J}(\mathbf{W})}{\partial\mathbf{W}}=-\mathbf{X}^{\mathsf{T}}\mathbf{XW}$
The matrix $\mathbf{W}$ in the above relation is that available in the previous iteration: $\mathbf{W}^{(t-1)}$. Then, the gradient updating equation is: $(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})\mathbf{W}^{(t-1)}$. Therefore, at iteration $t=1,2,\ldots$ and given column orthogonal matrix $\mathbf{W}^{(t-1)}$, the gradient descent optimization search method consists of two alternate steps to obtain the update:
$\tilde{\mathbf{W}}$ $=$ $\mathbf{G}\mathbf{W}^{(t-1)}$ (20)
$\mathbf{W}^{(t)}$ $=$ $Gram-Schmidt(\tilde{\mathbf{W}})$ (21)
where $\mathbf{G}\triangleq(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$, $\eta$ is a positive step parameter and $q\geq 1$ is an integer. The settings for $\eta$ and $q$ are discussed later. The method to use for obtaining the column orthogonal matrix $\mathbf{W}$ is summarized in algorithm 1.
Data: matrix $\mathbf{X}$ of size $m\times n$, ($m\geq n$), settings $\eta$, $q$ and $\epsilon$;
Result: matrix $\mathbf{W}$ with orthogonal columns;
1 Initialization: select randomly entries of a matrix $\mathbf{W}$ from a normal or an uniform distribution and perform orthogonalization $\mathbf{W}^{(0)}=$ Gram-Schmidt($\mathbf{W}$), compute $\mathbf{G}=(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$;
2 for $t\leftarrow 1,2,\ldots$ do
3       compute $\tilde{\mathbf{W}}$ using relation (20);
4       compute $\mathbf{W}^{(t)}$ using relation (21);
5       $\delta=\|\mathbf{W}^{(t)}-\mathbf{W}^{(t-1)}\|_{\mathsf{F}}^{2}$;
6       if $\delta\leq\epsilon$ then
7            stop;
8       end if
9
10 end for
Algorithm 1 Gradient search for columns orthonormal matrix $\mathbf{W}$
This algorithm is similar to the power method, see below.
2.1.1 Factors for the eigendecomposition of $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and for the SVD of $\mathbf{X}$
The algorithm 1 leads to a column orthogonal matrix $\mathbf{W}$ that fits the matrix $\mathbf{X}$. Using relation (8), the eigenvalues of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ are the entries of the diagonal matrix $\boldsymbol{\Lambda}$. These eigenvalues are the diagonal entries of the matrix $(\mathbf{XW})^{\mathsf{T}}\mathbf{XW}$ which is symmetric. The diagonal matrix of the SVD is: $\mathbf{D}=\boldsymbol{\Lambda}^{1/2}$. Finally the matrix $\mathbf{U}$ of the SVD is: $\mathbf{U}=\mathbf{XW}\boldsymbol{\Lambda}^{-1/2}$.
2.1.2 Relation with the power method
Algorithm 1 is in fact the power method allowing to obtain all eigenvalues together and not one by one as done using the classical power method algorithm 2, see Annexes. The matrix $\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)$ is used in this algorithm instead of the matrix $\mathbf{A}\triangleq\mathbf{X}^{\mathsf{T}}\mathbf{X}$. These two matrices have the same eigenvectors. Indeed, the eigenvector $\mathbf{z}$ associated with the eigenvalue $\gamma$ of the matrix $\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)$ allows to write:
$(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})\mathbf{z}=\gamma\mathbf{z}\Rightarrow\mathbf{X}^{\mathsf{T}}\mathbf{Xz}=\left(\frac{\gamma-1}{\eta}\right)\mathbf{z}$
The above relation means the matrices $\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)$ and $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ have the same eigenvectors which are associated with related eigenvalues, $j=1,2,\ldots,n$:
$\lambda_{j}$ $=$ $(\gamma_{j}-1)/\eta$ (22)
$\gamma_{j}$ $=$ $\eta\lambda_{j}+1$ (23)
where $\lambda_{j}$ and $\gamma_{j}$ are the eigenvalues of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$, respectively. Relation (23) shows that the smallest eigenvalue of the matrix $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$ is one, meaning this matrix is always of full rank even for rank deficient matrix $\mathbf{X}$. Safety computation of the norm $\|\mathbf{W}^{(t)}-\mathbf{W}^{(t-1)}\|_{\mathsf{F}}^{2}$, step 5 of alogithm 1, at any iteration requires to have the same number of columns for the matrix $\mathbf{W}$. When the matrix $\mathbf{X}$ is not of full rank, there exit a matrix $\mathbf{W}$ of size $n\times r$, $r<n$, allowing to have an exact reconstruction of the $\mathbf{X}$. Therefore, using alone the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ in algorithm 1 may lead a problem for computing the above norm if we are not careful.
2.1.3 Convergence analysis and choice of the settings $\eta$ and $q$
From the convex property of the objective function, the algorithm 1 is convergent. The rate of convergence depends especially on the behavior of the singular values of the matrix $\mathbf{X}$ and on the precision ($\epsilon$) required. The rate of convergence for the probabilistic SVD is analyzed in [17, 16] where faster convergence is observed for matrix which singular values decay quickly. That means, larger the difference between successive singular values is, faster is the convergence. This result corresponds to the modulus conditions for the power method. Relation (13) shows a possibility to modify the singular values decay using power $q$ of the matrix in the eigendecomposition. From relation (23), a small value (near zero) for the parameter $\eta$ leads to have all eigenvalues for matrix $\mathbf{G}$ close to one. On the other hand, a value greater to $1$ for the parameter $\eta$ leads to increasing the eigenvalues of the matrix $\mathbf{G}$. Therefore, the choices $q>1$ for the power parameter and $\eta\geq 1$ for the step parameter, likely lead to rapid singular values decay for the matrix $\mathbf{G}$.
The computation of the matrix $\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)^{q}$ may lead to numerical problem when $\mathbf{X}$ has some entries with very small or high values. This square matrix of order $n$ is computed once before starting the iterative search. At each iteration, a matrix multiplication and a Gram-Schmidt (GS) orthogonalization are performed. The modified GS orthogonalization method is summarized in algorithm 3, in the Annexes. The matrices multiplication and the GS orthogonalization are the time-consuming part of the proposed method which is implemented in the psvd R package of the R environment, [35] (https://cran.r-project.org/web/packages/psvd/index.html). A C code is used to speed-up the matrices multiplication and the GS orthogonalization calculator.
2.1.4 Small size matrice illustrative examples
To illustrate the choice of settings, three academic example matrices, relation (24), and the Fisher’s iris data, [11], were used. The iris dataset corresponds to measurements of the length and the width of the sepal and the petal for $3$ iris species: setosa, versicolor and virginica. $50$ measurements were observed for each species leading to $m=150$ and $n=4$.
$X_{a}=\left(\begin{array}[]{ccc}1&1&1\\ 0&2&1\\ 1&0&1\end{array}\right)\mbox{ ; }X_{b}=\left(\begin{array}[]{cccc}3&1&9&2\\ 10&4,&8&6\\ 7&6&12&1\\ 11&2&5&9\\ 1&1&1&0\end{array}\right)\mbox{ ; }X_{c}=\left(\begin{array}[]{ccccc}22&10&2&3&7\\ 14&7&10&0&8\\ -1&13&-1&-11&3\\ -3&-2&13&-2&4\\ 9&8&1&-2&4\\ 9&1&-7&5&-1\\ 2&-6&6&5&1\\ 4&5&0&-2&2\end{array}\right)$ (24)
The singular values associates with these examples are shown in Table 1. The second and the third eigenvalues for the matrix $\mathbf{X}_{c}$ are very close and the last two eigenvalues are zero.
Table 1: Singular values
# $X_{a}$ $X_{b}$ $X_{c}$ Iris
1 2.80193774 26.02508484 35.32704347 95.95991387
2 1.44504187 9.31733797 20 17.76103366
3 0.24697960 3.29881377 19.59591794 3.46093093
4 - 0 0 1.88482630
5 - - 0 -
Table 2: Convergence iterations
$q=1$
$\eta$ 0.01 0.1 1 10 100
$X_{a}$ 500 81 20 16 16
$X_{b}$ 156 25 11 11 11
$X_{c}$ 420 345 342 341 341
Iris 231 42 21 18 18
$q=2$
$\eta$ 0.01 0.1 1 10 100
$X_{a}$ 328 43 11 9 9
$X_{b}$ 82 14 6 6 6
$X_{c}$ 221 184 180 179 184
iris 120 22 11 10 10
For fixed power parameter $q=1,2$, the step parameter $\eta$ was varied. The number of iterations required for $8$ digits convergence to the values in Table 1 are shown in Table 2. Using $q=3$ or $q=4$ necessitates the same number of operations in the initialization step. For some test examples not shown, using $q=4$ do not show a significant reduce of the number of iterations than using $q=2$. In addition, using $q=4$ may lead to huge entries in the matrix $(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$ for some examples. Therefore, the choice $q=2$ is a good compromise.
2.1.5 High-dimensional example: trachea data
The trachea dataset corresponds to gene expression profiling using the sequencing technology. The data was downloaded from the Gene Expression Omnibus, [3], using the access number GSE103354, file: GSE103354_Trachea_fullLength_TPM.txt.gz. This dataset consists of full-length scRNA-seq for $301$ trachea segments of C57BL/6 wild-type mice, [29]. A total of $23,420$ genes (transcripts) are reported. There are $4,836$ transcripts which have zero count values for the $301$ cells. These transcripts were removed and the remaining dataset has $18,584$ rows.
Figure 1: Log scale eigenvalues of the trachea dataset
Figure 1 shows the behavior of the $\log$ scale singular values. For this example, the first $87$ singular values capture $90\%$ of information and the first $159$ singular values capture $95\%$ of information. One can observe a fast decay for the first $20$ singular values and a slow decay for the others.
Five methods available in the literature were used to show the performance of the proposed method. The calculations were performed on the same laptop computer equipped with i5-1135G7 processor, 8 GB of RAM, under Microsoft Windows 11 family and R environment version 4.4.1, [35]. The psvd R package of the proposed method is available from the comprehensive R archive network, http://cran.r-project.org/web/packages/psvd/. The classical SVD method is available through the R build-in svd() function which computes all singular values. The Lanczos algorithm based method for computing all singular values is available in the svd R package, [21]. A C++ library for large scale eigenvalue problem is available in the Spectra tool, https://github.com/yixuan/spectra. The RSpectra R package allows to use this library for computing the first $r$ singular values, [34]. The irlba R package, [2], also allows to compute the first $r$ singular values. The probabilistic SVD method is implemented in the rsvd R package, [10]. When possible, the largest $20$, $50$, $100$ or $150$ singular values were computed. Each method was run $10$ times using default settings and the average results were reported. For the proposed method, the following settings were used: $q=2$, tolerance ϵ=italic-$\epsilon=$1e-8 and $\eta=10$. The classical svd() function was used as reference and three parameters were examined: the runtime, percentage rate of reconstruction and the means squared error (MSE).
Table 3: Comparison results
Method runtime (s) reconst. Rate % MSE
$r=20$ svd 3.77 80.83 -
propack.svd 4.58 80.83 1.39e-8
RSpectra 0.53 80.83 1.53e-8
irlba 0.42 80.83 0.00160
rsvd 1.04 80.83 320.106
psvd 1.66 80.83 2.35e-8
$r=50$ svd 3.77 86.21 -
propack.svd 4.58 86.21 1.39e-8
RSpectra 1.29 86.21 1.56e-8
irlba 1.36 86.21 0.00554
rsvd 2.42 86.15 1740.22
psvd 2.61 86.21 7.77496
$r=100$ svd 3.77 91.11 -
propack.svd 4.58 91.11 1.39e-8
RSpectra 2.12 91.11 0.69e-8
irlba 5.42 91.11 0.00989
rsvd 4.92 91.01 2162.08
psvd 4.50 91.11 38.2471
$r=150$ svd 3.77 94.47 -
propack.svd 4.58 94.47 1.39e-8
RSpectra 3.18 94.47 1.51e-8
irlba 14.53 94.47 0.01732
rsvd 7.54 94.38 1688.06
psvd 6.93 94.47 46.3998
The results obtained are summarized in Table 3 where the proposed method appears in bold.
Runtime.
All singular values are computed togheter using the classical svd and the propack.svd methods. The former method is slightly faster than the later. When the first $20$ are required, the RSpectra, irlba, rsvd and psvd methods use less runtime compare to the svd and propack.svd methods. For this example, when $r=150$, the RSpectra is the fastest followed by classical svd, propack.svd, psvd, rsvd and irlba. The runtime of this later method grows quickly when the number of the required singular values increases.
Reconstruction rate.
Relation (3) was used to compute the reconstruction rate. The denominator is the same for all method and its value is obtained using the classical svd method results. The numerator is computed using the results of each method. The reconstruction rate is relatively the same for all methods. A very small difference appears for the rsvd method when $r\geq 50$.
Mean squared error.
The differences between the singular values of each method and those obtained using the classical method where computed and used to calculate a mean squared error (MSE). The propack.svd and the RSpectra methods have quasi similar reconstruction rate as the classical svd method. The irlba method has near zero MSE value. The methods psvd and rsvd have more than zero MSE values. The psvd method MSE can be improved by using small tolerance $\epsilon$ value. The rsvd method has relatively higher MSE value and can lead to a significant gap between singular values, see Table 4 showing the last $10$ singular values for $r=100$. From the results of the rsvd method, we observed that the first $50$ estimates obtained using $r=100$ are more accurate than those obtained using $r=50$. The first $20$ singular values obtained by all methods are relatively accurate.
Table 4: $91$-th to $100$-th singular values
i svd psvd rsvd
91 10847.51 10846.65 10390.96
92 10804.19 10805.04 10331.54
93 10638.48 10638.33 10260.26
94 10600.40 10600.56 10040.75
95 10413.53 10413.28 9963.21
96 10376.98 10377.18 9920.82
97 10332.29 10332.33 9820.13
98 10289.21 10289.21 9691.56
99 10195.62 10195.62 9639.37
100 10002.21 10002.21 9432.25
The irlba and the rsvd methods cannot be used for small size matrices. Therefore, only RSpectra and the proposed psvd methods can be used to compute the $r$ largest singular values for any matrix.
3 Applications
In the above presentation, the eigendecomposition and the SVD consist of linear algebra manipulations of the matrix $\mathbf{X}$ entries. Truncating of the sum in (2) to $r<n$ leads to a $r$ rank approximation of the matrix $\mathbf{X}$. A direct application of the $r$ largest eigenvectors associated with the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ is the principal component analysis.
3.1 Principal component analysis
In the SVD analysis, no assumption was used on the statistical properties for the entries of the matrix $\mathbf{X}$. In practical applications, the row $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, of this matrix is a measurement sample of a phenomenon. Let us assume that the $n$-dimensional vector $\mathbf{x}$ is an observation of a random vector $\mathcal{X}$ which entries are the variables of the phenemenon under study. The mean and the covariance of this random vector are: $\mathbb{E}[\mathcal{X}]\triangleq\boldsymbol{\mu}$ and $\mathbb{V}[\mathcal{X}]=\mathbb{E}[(\mathcal{X}-\boldsymbol{\mu})(\mathcal{X}-\boldsymbol{\mu})^{T}]\triangleq\boldsymbol{\Sigma}$. The diagonal entries of the matrix $\boldsymbol{\Sigma}$ are the variances of the $\mathcal{X}$ entries: $cov(\mathbf{x}_{j},\mathbf{x}_{j})=\sigma_{j}^{2}$, $j=1,2,\ldots,n$. The principal component analysis (PCA) consists of the eigendecomposition of the covariance matrix, followed by a representation of the observations and variables in the reduced space axes, [19, 20]. Very often, the random vector $\mathcal{X}$ is assumed to be of zero mean. When the entries of $\mathcal{X}$ are heterogenous, there can be a large gap between the amplitudes of the variances $\sigma_{j}^{2}$, $j=1,2,\ldots,n$. In this case the role of the variables with small variances is hidden in the interpretation of the PCA results. To have homogeneous variance data, the centered value of each entry should be divided by the corresponding standard deviation $\sigma_{j}$. The covariance matrix becomes a correlation matrix which has only ones as diagonal entries.
3.1.1 PCA for finite number of observations
When $m$ observations or objects are available, only estimates can be obtained for the mean vector $\boldsymbol{\mu}$ and the covariance matrix $\boldsymbol{\Sigma}$. The sample mean $\bar{\mathbf{x}}$ and the sample variance $s_{j}^{2}$ of the $j$-th entry of the observations $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, are estimated as:
$\bar{\mathbf{x}}=\frac{1}{m}\sum_{i=1}^{m}\mathbf{x}_{i}\mbox{ ; }s_{j}^{2}=\frac{1}{m-1}\sum_{i=1}^{m}(x_{ij}-\bar{x}_{j})^{2}\mbox{ , }j=1,2,\ldots,n$
From the entries $x_{ij}$ of the matrix $\mathbf{X}$, the entries $x_{ij}^{c}$ of the centered matrix $\mathbf{X}_{c}$ and the entries $x_{ij}^{s}$ of the standardized matrix $\mathbf{X}_{s}$ are obtained as ($i=1,2,\ldots,m$, $j=1,2,\ldots,n$):
$\mathbf{X}_{c}$ ::: $x^{c}_{ij}=x_{ij}-\bar{x}_{j}$
$\mathbf{X}_{s}$ ::: $x^{s}_{ij}=(x_{ij}-\bar{x}_{j})/s_{j}$
Estimates for the sample covariance and the sample variance matrices are obtained using $\mathbf{X}_{c}^{\mathsf{T}}\mathbf{X}_{c}/m$ and $\mathbf{X}_{s}^{\mathsf{T}}\mathbf{X}_{s}/m$, respectively. Using the expressions between the SVD factors and those of the eigendecomposition, the PCA result can be obtained from the SVD results of the matrix $\mathbf{X}_{c}$ or the matrix $\mathbf{X}_{s}$.
3.1.2 Representation of objects or variables in reduced space
The matrices $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and $\mathbf{XX}^{\mathsf{T}}$ have the same eigenvalues and related eigenvectors. The eigenvalue $\lambda$ allows to write:
$\mathbf{X}^{\mathsf{T}}\mathbf{Xv}$ $=$ $\lambda\mathbf{v}$ (25)
$\mathbf{XX}^{\mathsf{T}}\mathbf{u}$ $=$ $\lambda\mathbf{u}$ (26)
where $\mathbf{v}$ and $\mathbf{u}$ are the eigenvectors. Relation (25) allows to write: $\mathbf{XX}^{\mathsf{T}}\mathbf{Xv}=\lambda\mathbf{Xv}$, meaning that $\mathbf{u}=\mathbf{Xv}$. Relation (26) can be used in a similar way to show that $\mathbf{v}=\mathbf{X}^{\mathsf{T}}\mathbf{u}$. Let $r$ be the reduced dimension used in the PCA, typically $r=2$ or $3$ for a visual representation. From the PCA results of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$, expressions for the objects in the $r$-dimension variable space are $\mathbf{X}\mathbf{V}_{r}$, where $\mathbf{V}_{r}$ is a matrix with the $r$ largest eigenvectors. Using the expression for the matrix $\mathbf{U}$ in relation (14), expressions for the variables in the $m$-dimensional space are obtained as: $\mathbf{X}^{\mathsf{T}}\mathbf{XV}_{r}\boldsymbol{\Lambda}_{r}^{-1/2}$, where $\boldsymbol{\Lambda}_{r}^{-1/2}$ is the inverse of the diagonal matrix obtained using the $r$ largest nonzero eigenvalues.
The PCA results plot for the iris data is shown in Figure 2.
Figure 2: Plot of the iris observations in the first and third principal components axes.
3.2 Autoencoder
An autoencoder (AE) is a neural network (NN) model which allows to encode input, store the result internally, and finally decode the stored result to obtain an output. The output is not necessary strictly equal to the input. In comparison, the SVD leads to the matrices $\mathbf{U}$, $\mathbf{D}$ and $\mathbf{V}$ which are an internal representation of the input matrix $\mathbf{X}$. Then, using a rank $r$ value, an approximation of the input is obtained: $\hat{\mathbf{X}}=\mathbf{U}_{r}\mathbf{D}_{r}\mathbf{W}_{r}^{\mathsf{T}}$. This approximation matrix is of rank $r$ and the SVD uses linear operations to solve the dimensionality reduction problem or data compression. The data can also be compressed using a feedforward NN model with few layers, [30, 4, 22]. A NN layer consists of a weighted combination of an input, $\mathbf{x}$, followed by a transformation function $\varphi(.)$ to obtain an output: $\mathbf{u}=\mathbf{x}^{\mathsf{T}}\mathbf{W}+\mathbf{w}_{o}$; $\mathbf{v}=\varphi(\mathbf{u})$, where $\mathbf{W}$ is weight matrix and $\mathbf{w}_{o}$ is a threshold or bias vector. The transformation operates on each entry of the vector $\mathbf{u}$. Three examples of the transformation function are the sigmoid, $\sigma(u)=1/(1+e^{-u})$, the rectified linear unit, $ReLU(u)=max(0,u)$, and the identity $I(u)=u$. A two layers feedforward NN model has been used in [22] to perform a nonlinear SVD for obtaining an approximation of the input $\mathbf{X}$. Each row of the matrix $\mathbf{X}$ is an $n$-dimensional vector $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, and the output of the encoding layer is obtained as:
$\mathbf{v}_{i}=\sigma\left(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{W}_{e}+\mathbf{w}_{e}\right)$
where $\mathbf{W}_{e}$ is a weight matrix of size $n\times r$ and $\mathbf{w}_{e}$ is a bias vector of order $n$. The output of the decoding layer is obtained as:
$\hat{\mathbf{x}}_{i}=\sigma\left(\mathbf{v}_{i}^{\mathsf{T}}\mathbf{W}_{d}+\mathbf{w}_{d}\right)=\sigma\left(\sigma\left(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{W}_{e}+\mathbf{w}_{e}\right)^{\mathsf{T}}\mathbf{W}_{d}+\mathbf{w}_{d}\right)$
where $\mathbf{W}_{d}$ is a weight matrix of size $r\times n$ and $\mathbf{w}_{d}$ is a bias vector of order $n$. The input $\mathbf{x}_{i}$ and its approximation by the decoder output are used to define an $\ell_{2}$ norm:
$\|\mathbf{x}_{i}-\hat{\mathbf{x}}_{i}\|$
From the $m$ observations, the weight matrices, $\mathbf{W}_{e}$ and $\mathbf{W}_{d}$, and the bias vectors, $\mathbf{w}_{e}$ and $\mathbf{w}_{d}$ are obtained after minimization of the $\ell_{2}$-norm using a gradient search optimization algorithm. Instead of using a single layer in the encoder, $k$ layers, $k>1$, can be used to progressively increase the compression rate. In this case, an inverse operation is done using $k$ decoding layers. The approximation matrix obtained using the feedforward NN model is not necessary of reduced rank even the number of columns of the encoder weight matrix is $r<n$. However, equivalent approximation to SVD result can be obtained using the AE model. The difference matrix $\mathbf{X}-\hat{\mathbf{X}}$ can be viewed as the noise part of the data. Therefore, the approximation $\hat{\mathbf{X}}$ can be used in subsequent supervised or unsupervised data analysis methods.
3.2.1 Identity transformation function
The identity transformation function has been used in [33, 36]. The transpose of the matrix $\mathbf{X}$ is used in [33] leading to $\|\mathbf{X}-\mathbf{WW}^{\mathsf{T}}\mathbf{X}\|$ while the constrained optimization objective function of the relation (16) is used in [36, Chapter 7]. In these two references, the classical gradient search optimization method is used to obtain the matrix $\mathbf{W}$. The algorithm 1 solves the same problem by converting the gradient search into a power method algorithm and adds a convergence acceleration through the choice of the parameter $q$. When an estimation is available for the weight matrix $\mathbf{W}$, then an proximation of the input matrix $\mathbf{X}$ is: $\hat{\mathbf{X}}=\mathbf{XWW}^{\mathsf{T}}$. Therefore, an approximation of the input vector $\mathbf{x}_{i}$ is: $\hat{\mathbf{x}}_{i}=\mathbf{x}_{i}^{\mathsf{T}}\mathbf{WW}^{\mathsf{T}}$.
MNIST data.
For illustration, the MNIST data, [26], was used. This dataset consists of $60,000$ training images and $10,000$ test images. Each image of size $28\times 28$ is an observation or vector of dimension $784$. The entries of this vector are pixels which values are from the set $\{0,1,\ldots,255\}$. The behavior of the singular values and the cumulative sum of the reconstruction percentage rate of the training images data are show in Figure 3. The last $72$ singular values are zero. The total information captured by the first $50$, $100$ or $256$ largest singular values are $45.41\%$, $60.45\%$ and $82.63\%$, respectively.
(a) log scale singular values
(b) cumulative reconstruction rate
Figure 3: Log scale singular values and cumulative percentage reconstruction rate
The largest $256$ singular values were computed using the training images data. The weight matrix obtained was used to estimate approximates for the first $5$ test images. The results are shown in Figure 4, where the first column is the test image and the $3$ following columns are for the approximation using the first $50$, $100$ or $256$ singular values, respectively. The compression rates for these approximations are $15.68$, $7.84$ and $3.06$, respectively.
Figure 4: First 5 test images, from left to right: input, use of the $50$, $100$ or $256$ largest singular values to obtain approximation.
4 Conclusion
A new method is proposed for computing the singular values of a matrix. This method allows to compute all or a subset of the largest singular values and associated vectors. It can be applied to small or high dimensional dataset. The accuracy of the results depends on the user tolerance setting. More accuracy may necessitate a long runtime depending on the behavior of the singular values decay. The proposed method especially consists of searching for an eigendecomposition of a square symmetric matrix by using a gradient method to minimize a Frobenius norm. The method was applied to principal component analysis and to neural network autoencoder model.
Annexes
The proposed method is implemented in the psvd R package available on comprehensive R archive network [35]: http://cran.r-project.org/web/packages/psvd/. In the first version, the power parameter $q$ is fixed to $2$ and only a sample from standard normal distribution is used as initial projection matrix. The basic power method algorithm is given below to show the similarity with the proposed algorithm. For completeness, a numerically stable Gram-Schmidt orthogonalization algorithm is provided below. The iterative part of the proposed algorithm is the more runtime consuming. A C code is used for this part in the psvd R package. Reader can use her/his preferred computing language to implement all these algorithms.
Power method
Data: square matrix $\mathbf{A}$
Result: orthonormal vector $\mathbf{v}$
1 Initialization: choose an initial vector $\mathbf{v}^{(0)}$ and compute $\mathbf{v}^{(0)}\leftarrow\mathbf{v}^{(0)}/\|\mathbf{v}^{(0)}\|$, set tolerance $\epsilon$;
2 for $t\leftarrow 1,2,\ldots$ do
3       Compute a vector $\mathbf{z}=\mathbf{Av}^{(t-1)}$;
4       Update vector: $\mathbf{v}^{(t)}=\mathbf{z}/\|\mathbf{z}\|$;
5       $\delta\leftarrow\|\mathbf{v}^{(t)}-\mathbf{v}^{(t-1)}\|$;
6       if $\delta\leq\epsilon$ then
7            stop;
8       end if
9
10 end for
Algorithm 2 Power method
.0.1 Modified Gram-Schmidt algorithm
Data: Input = $\mathbf{W}$ has $n$ columns and $m\geq n$ rows
Result: Output = $\mathbf{V}$, $\mathbf{v}_{j}$ is the $j$-column of $\mathbf{V}$
1 $\mathbf{V}\leftarrow\mathbf{W}$;
2 for $i\leftarrow 1$ to $n$ do
3       $r_{i}\leftarrow\left\langle\mathbf{v}_{i},\mathbf{v}_{i}\right\rangle^{1/2}$;
4       $\mathbf{v}_{i}\leftarrow\mathbf{v}_{i}/r_{i}$;
5       if $i<n$ then
6             for $j\leftarrow i+1$ to $n$ do
7                   $r_{j}\leftarrow\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle$;
8                   $\mathbf{v}_{j}\leftarrow\mathbf{v}_{j}-r_{j}\mathbf{v}_{i}$;
9
10             end for
11
12       end if
13
14 end for
Algorithm 3 Modified Gram-Schmidt orthogonalization method
where $\langle\mathbf{x},\mathbf{y}\rangle$ is the inner product of the vectors $\mathbf{x}$ and $\mathbf{y}$.

### v4

@@
 Abstract
-The singular value decomposition (SVD) allows to put a matrix as a product of three matrices: a matrix with the left singular vectors, a matrix with the positive-valued singular values and a matrix with the right singular vectors.
-There are two main approaches allowing to get the SVD result: the classical method and the randomized method.
-The analysis of the classical approach leads to accurate singular values.
-The randomized approach is especially used for high dimensional matrix and is based on the approximation accuracy without computing necessary all singular values.
-In this paper, the SVD computation is formalized as an optimization problem and a use of the gradient search algorithm.
-That results in a power method allowing to get all or the first largest singular values and their associated right vectors.
-In this iterative search, the accuracy on the singular values and the associated vector matrix depends on the user settings.
-Two applications of the SVD are the principal component analysis and the autoencoder used in the neural network models.
+The singular value decomposition (SVD) allows to write a matrix as a product of a left singular vectors matrix, a nonnegative singular values diagonal matrix and a right singular vectors matrix.
+Among the applications of the SVD are the principal component analysis, the low-rank matrix approximation and the solving of a linear system of equations.
+The methods used for computing this decomposition allow to get the complete or partial result.
+For very large size matrix, the probabilistic methods allow to get partial result by using less computational load.
+A power method is proposed in this paper for computing all or the $k$ first largest SVD subspaces for a real-valued matrix.
+The $k$ first right singular vectors of this method are the $k$ columns of a neural network encoder weight matrix.
+The accuracy of this iterative search method depends on the behavior of the singular values and the settings of the gradient search optimizer used.
+A R package implementing the proposed method is available at https://cran.r-project.org/web/packages/psvd/index.html.
 Keywords:
-eigendecomposition, power method, singular values, principal components, autoencoder, neural network.
+eigen values, singular vectors, singular values, power method, principal components, autoencoder, neural networks.
 1 Introduction
-Nowadays, we are dealing with massive data in our current activities: words count in text documents, image, gene expression from sequencing analyses, etc.
-Such data are typically organized in a matrix of size $m\times n$ where $m$ and $n$ are respectively the number of rows and columns of this matrix.
-The values $m$ and $n$ also define the matrix: $\min(m,n)$.
-We assume that $m\geq n$, this condition can be met by using the matrix transpose.
-One characteristic of the matrix associated with the words count in a text document, image, single-cell RNA sequencing or similar data, is the number of zero entries which can be high, up to $50\%$ or more.
-The singular value decomposition (SVD) allows to put a matrix into a product of three matrices.
-The properties of these factor matrices allow to write the matrix as a sum of rank one matrices, [9].
-The contribution of each term in this sum is quantified by a nonnegative value called singular value.
-The classical SVD computation methods proceed in two steps: bidiagonalization or rank-revealing decomposition of the matrix, then, search for the factor matrices, [13, 14, 6, 8].
-Usually, all singular values and associated vectors are computed using a classical method.
-In some data analysis methods, the smallest singular value and corresponding vectors may be important, [8].
-For the various high dimensional data encountered todays, compact form of the data is often required for storing purpose or for revealing some links between data objects.
-That is why randomized methods for computing SVD come in play.
-Indeed, the first important singular values are necessary to get some useful information from the data.
-These methods also proceed into two steps: search for projection feature space, then obtention of the SVD factor matrices, [17, 16].
-In this paper, a method for computing SVD is proposed.
-This algorithm is especially suitable for searching for few singular values of a high dimensional data but it also applied to small size matrices.
-The proposed algorithm is inspired from the probabilistic random search approaches by [17, 16].
-The difference comes from our formulation of the SVD computing as an objective function optimization problem and the use of a power method to get the projection matrix.
-1.1 Classical SVD computation
-The SVD allows to decompose any matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$ as a product of three other matrices, [9, 28, 15, 18]:
+Nowadays, we are dealing with massive data in our current activities: words count in text documents, images, gene expression from sequencing analyses, etc.
+Such data are typically organized in a matrix of size $m\times n$ where $m$ and $n$ are the number of rows and columns, respectively.
+We assume that $m\geq n$, this condition can be met by using a matrix transpose.
+A compact form of a large size data is often required for storing purpose or for revealing some links between the samples.
+The singular value decomposition (SVD) allows to express a real-valued matrix as a product of a left orthogonal, a diagonal and a right orthogonal matrices, [14, 58].
+The entries of the diagonal matrix are the nonnegative-valued singular values and the columns of the orthogonal matrices are the singular vectors.
+The properties of these vectors allow to write the matrix as a sum of rank one matrices.
+The contribution of each term in this sum depends on the weight of the associated singular value.
+One strategy for obtaining the SVD result consists of computing a symmetric matrix followed by some transformations.
+In a first category of transformation methods, a bidiagonalization or rank revealing decomposition is done followed by a search for the factor matrices [21, 22, 10, 13, 12].
+All singular values and singular vectors are computed at a time using this category of methods.
+The quantification in SVD perturbation uncertainties is addressed in [63] while the convex set-oriented SVD method proposed in [16] overcomes the limitation of the nominal SVD method in the presence of uncertainties.
+In a second category of transformation methods, an iterative search allows to get the factor matrices entries [36, 28, 1, 56, 33, 65].
+This second category of methods allows to compute the $k$ largest or smallest singular values and singular vectors.
+Another strategy for computing the SVD consists of a direct algebraic manipulation of the matrix entries.
+The refinement method proposed in [48] is based on matrices multiplication for computing the SVD of a full columns rank matrix.
+The randomized methods allow a direct search for an orthogonal matrix associated with the first $k$ largest singular values, [19, 26, 25].
+When the data matrix size is large, the first $k$ largest subspaces are often enough to have a good data matrix approximation.
+However, to have accurate $k$ first singular values using the randomized method, a matrix of order greater $k$ should be used, see Table 4.
+The neural network (NN) model is nowdays used in many applications, [6, 39, 64, 41].
+This model has a weight matrix which entries are computed from a training dataset.
+The NN autoencoder model consists of two functions (encoding and decoding) allowing to store internally the input and then output it in a desirable format.
+This model has numerous applications: linear algebra, data denoising, regression analysis, texts and images generation, natural language processing, supervised and unsupervised classification.
+For the unsupervised classification or data clustering, the autoencoder model allows to get lower dimensional or denoised data which are used to obtain clusters.
+The singular values and the singular vectors of the SVD are the encoding results, while the decoding can consist of using a reconstruction form the $k$ first singular values and singular vectors.
+A deep embedding clustering (DEC) method is proposed in [61] by using an autoencoder to have a lower dimensional dataset, the decoder output.
+The cluster centers in DEC are learned from this lower dimensional data after a minimization of an objection function involving a Kullback-Leibler divergence and the t-distribution.
+The deep divergence clustering method proposed in [32] uses convolutional layers, autoencoder, for images datasets.
+An autoencoder is used in the deep-k-means method proposed in [17].
+A dynamic autoencoder is used in [47] for clustering a dataset by gradually and smoothly eliminating the reconstruction objective function in favor of a construction one.
+The autoencoder model parameters and the clustering parameters are simultaneously learned in the method proposed in [7] by using a certain class of gaussian mixtures based objective function.
+The deep convolutional embedded clustering algorithm proposed in [43] uses a convolutional autoencoder followed by convolutional neural network.
+The unsupervised companion objectives method proposed in [60] allows to reduce the mismatch which occurs when the optimization of an objective (the autoencoder reconstruction loss function) has a negative impact on the optimization of another objective (the clustering loss function).
+In data matrix approximation applications, it has been shown that the singular vectors can be the NN model weight matrix columns, [49, 8, 4, 35, 62].
+An independent optimization method is very often used to get the singular vectors.
+For the iterative search method proposed, the optimizer output singular vectors are the NN weight matrix columns.
+The right singular vectors matrix of the probabilistic method correspond to an NN weight matrix.
+However, the probabilistic method results are less accurate when the number of the singular values increases.
+The proposed algorithm is inspired from the probabilistic random search approaches, [26, 25].
+The difference comes from a minimization of an objective function and the use of a power method to get simultaneously the singular vectors.
+The proposed method is more accurate, can be used for any matrix and also allows to compute the complete SVD result.
+1.1 Computing the SVD
+The SVD of any matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$ is, [44, 24, 29]:
 $\mathbf{X}=\mathbf{UD}\mathbf{W}^{\mathsf{T}}$ (1)
-where $\mathbf{U}$ and $\mathbf{W}$ are square column orthogonal matrices of order $m$ and $n$, respectively, $\mathbf{D}$ is a matrix of size $m\times n$.
-Only the diagonal entries of the upper square part of the matrix $\mathbf{D}$ are nonzero and are the positive-valued singular values: $d_{11}\geq d_{22}\geq\ldots\geq d_{nn}$.
-In the thin SVD, the matrix $\mathbf{U}$ is of size $m\times n$ and $\mathbf{D}$ is a square diagonal matrix of order $n$.
-Let us observe that $-\mathbf{U}$ and $-\mathbf{W}$ are jointly also solution of the singular value decomposition in relation (1).
-The column orthogonal property of the matrices $\mathbf{U}$ and $\mathbf{W}$ allows to write: $\mathbf{U}^{\mathsf{T}}\mathbf{U}=\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{n}$, where $\mathbf{I}_{n}$ is the identity matrix of order $n$.
-Using a sequence of Householder transformations, a bidiagonal form is obtained: $\mathbf{X}=\mathbf{PGDH}^{\mathsf{T}}\mathbf{Q}^{\mathsf{T}}=\mathbf{PBQ}^{\mathsf{T}}$, where the upper square part of the matrix $\mathbf{B}$ is triangular with the same singular values as the matrix $\mathbf{X}$, $\mathbf{U}=\mathbf{PG}$ and $\mathbf{W}=\mathbf{QH}$.
-The singular values are easily computed from the matrix $\mathbf{B}$ leading to the matrices $\mathbf{U}$ and $\mathbf{W}$.
-The LDU (lower triangular, diagonal and upper triangular) decomposition is used by [8] to obtain the SVD.
+where $\mathbf{U}$ and $\mathbf{W}$ are the left and right orthogonal matrices of order $m$ and $n$, respectively, $\mathbf{D}$ is a matrix of size $m\times n$.
+The diagonal entries of the upper square part of the matrix $\mathbf{D}$ are nonnegative, the singular values: $d_{11}\geq d_{22}\geq\ldots\geq d_{nn}$.
+In the thin SVD, the matrix $\mathbf{U}$ is of size $m\times n$ and $\mathbf{D}$ is of order $n$.
+Let us observe that $\mathbf{U}$ and $\mathbf{W}$ are jointly also solution of the singular value decomposition in relation (1).
+The column orthogonal property of the matrices $\mathbf{U}$ and $\mathbf{W}$ allows to write: $\mathbf{U}^{\mathsf{T}}\mathbf{U}=\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{n}$, where $\mathbf{I}_{n}$ is a $n$-order identity matrix.
+Using the orthogonal property, relation (1) can be written as:
+$\mathbf{X}=\sum_{j=1}^{n}d_{jj}\mathbf{u}_{j}\mathbf{w}_{j}^{\mathsf{T}}$ (2)
+where the unit $\ell_{2}$-norm left and right singular vectors $\mathbf{u}_{j}$ and $\mathbf{w}_{j}$ are the $j$-th columns of the matrices $\mathbf{U}$ and $\mathbf{W}$, respectively, $\mathbf{u}_{k}^{\mathsf{T}}\mathbf{u}_{j}=1$, if $k=j$, 0 otherwise, idem for vectors $\mathbf{w}_{j}$.
+Again, for column $j$, the couples of vectors ($\mathbf{u}_{j},\mathbf{w}_{j}$) and ($\mathbf{u}_{j},-\mathbf{w}_{j}$) have the same contribution.
+If there are $n$ nonzero singular values $d_{jj}$, then the matrix $\mathbf{X}$ is of full rank.
+Otherwise, the number of nonzero scalars $d_{jj}$ is the actual rank of $\mathbf{X}$.
+The sum in relation (2) is often truncated to $r$ terms ($r<n$) leading to a lower rank matrix approximation.
+The percentage of the information associated with the $r$ first largest singular values is:
+$reconst.(\%)=100\sum_{j=1}^{r}d_{jj}\left/\sum_{j=1}^{n}d_{jj}\right$. (3)
+where the parameter $r$ is chosen to have the desired reconstruction rate.
+1.2 Eigenvalue decomposition
+The eigenvalue decomposition (EVD) or eigendecomposition can be performed for any square matrix, [23] . The EVD of the symmetric matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ is:
+$\mathbf{X}^{\mathsf{T}}\mathbf{X}=\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{\mathsf{T}}$ (4)
+where $\boldsymbol{\Lambda}$ is a diagonal matrix which entries are nonnegative real-valued scalars, the eigenvalues: $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}\geq 0$, $\mathbf{V}$ is an orthogonal matrix, $\mathbf{V}^{\mathsf{T}}\mathbf{V}=\mathbf{I}_{n}$.
+Like in relation (2), the EVD can be expressed as the sum of rank one matrices, the spectral decomposition:
+$\mathbf{X}^{\mathsf{T}}\mathbf{X}=\sum_{j=1}^{n}\lambda_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{\mathsf{T}}$ (5)
+where $\mathbf{v}_{j}$ is the $j$-th column of $\mathbf{V}$, a unit $\ell_{2}$-norm vector.
+The above sum can also be truncated to get a $r<n$ low-rank matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ approximation.
+In the principal component analysis, see below, the columns of the matrix $\mathbf{V}$ correspond to the projection subspaces.
+One can observe that $\mathbf{V}$ is also solution of the EVD, relation (4).
+This observation allows to fulfil the Perron-Frobenius theorem, [50, 20, 45], which stipulates that the entries of the main eigenvector of a nonnegative-valued matrix are all positive or zero.
+Nonnegative-valued matrices occur for many real-life data, e.g.
+genes expression and images.
+The main eigenvector is used in the ranking applications, [9, 37].
+To compute the eigenvalue with maximum modulus and the associated eigenvector, the iterative power method, [24], is very often used, see algorithm 2 in the Annexes.
+The convergence rate of the power method depends on the modulus of the second eigenvalue compared to the first, [24, 11].
+1.3 Singular value decomposition and eigenvalue decomposition
+It has been shown that the SVD factor matrices of a symmetric matrix are related to those of the EVD [21, 27, 58]. Indeed, using the expression (1) of $\mathbf{X}$ in the relation (4) allows to write:
+$\mathbf{X}^{\mathsf{T}}\mathbf{X}=(\mathbf{UDW}^{\mathsf{T}})^{\mathsf{T}}\mathbf{UDW}^{\mathsf{T}}=\mathbf{WD}^{2}\mathbf{W}^{\mathsf{T}}$
+Comparing the above expression to relation (4), the SVD factors using the EVD factors are:
+$\mathbf{W}=\mathbf{V}\mbox{ ; }\mathbf{D}=\boldsymbol{\Lambda}^{1/2}\mbox{ and }\mathbf{U}=\mathbf{XW}\boldsymbol{\Lambda}^{-1/2}$ (6)
+The right singular vectors of $\mathbf{X}$ are the eigenvectors of $\mathbf{X}^{\mathsf{T}}\mathbf{X}$.
+The entries of the diagonal matrix $\boldsymbol{\Lambda}$, the eigenvalues, are the diagonal entries of the symmetric matrix $(\mathbf{XW})^{\mathsf{T}}\mathbf{XW}$.
+1.4 Bidiagonalization methods for computing the SVD
+The SVD is computed in [24] by using a bidiagonalization of the matriox $\mathbf{X}$ or a QR decomposition of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$.
+In the first case, a sequence of Householder transformations leads to an upper bidiagonal matrix $\mathbf{B}$ and two other matrices, $\mathbf{U}_{1}$ and $\mathbf{V}_{1}$, [24, page 252].
+In the second case, the symmetric matrix allows to get a column orthogonal matrix $\mathbf{Q}$ and an upper triangular matrix $\mathbf{R}$, [24, page 448].
+In both cases, another manipulation of the factor matrices is used to get the SVD.
+The lower triangular, diagonal and upper triangular (LDU) decomposition is used in [12] to obtain the SVD.
 The diagonal entries of the triangular matrices $\mathbf{L}$ and $\mathbf{U}$ are all $1$.
 The diagonalization of the matrix $\mathbf{X}$ before obtaining the SVD leads to relatively accurate singular values.
-All singular values and their associated vectors are recovered using the classical SVD method.
-Using the column orthogonal property of the matrices $\mathbf{U}$ and $\mathbf{W}$, the decomposition in relation (1) can be written as:
-$\mathbf{X}=\sum_{j=1}^{n}d_{jj}\mathbf{u}_{j}\mathbf{w}_{j}^{\mathsf{T}}$ (2)
-where $d_{jj}$ are the singular values: $d_{11}\geq d_{22}\geq\ldots\geq d_{nn}$, the vectors $\mathbf{u}_{j}$ and $\mathbf{w}_{j}$ are the $j$-th columns of the matrices $\mathbf{U}$ and $\mathbf{W}$, respectively, $\mathbf{u}_{k}^{\mathsf{T}}\mathbf{u}_{j}=1$, if $k=j$, 00 otherwise, idem for vectors $\mathbf{w}_{j}$.
-Again, for column $j$, the couples of vectors ($\mathbf{u}_{j},\mathbf{w}_{j}$) and ($-\mathbf{u}_{j},-\mathbf{w}_{j}$) have the same contribution.
-If there are $n$ nonzero scalars $d_{jj}$, then the matrix $\mathbf{X}$ is of full rank.
-Otherwise, the number of nonzero scalars $d_{jj}$ is equal to the actual rank of the matrix $\mathbf{X}$.
-The sum relation (2) is often truncated to $r$ terms ($r<n$) leading to an approximate result of rank lower than data matrix.
-The percentage of the information captured by the $r$ largest singular values is quantified as:
-$reconst.(\%)=100\left(\sum_{j=1}^{r}d_{jj}\left/\sum_{j=1}^{n}d_{jj}\right.\right)$ (3)
-where the parameter $r$ is chosen to have the desired percentage in data reconstruction applications.
-The Frobenius norm of the matrix $\mathbf{X}$ is:
-$\|\mathbf{X}\|_{\mathsf{F}}^{2}=\sum_{i=1}^{m}\sum_{j=1}^{n}x_{ij}^{2}=\operatorname{tr}\left[\mathbf{X}^{\mathsf{T}}\mathbf{X}\right]=\operatorname{tr}\left[\mathbf{XX}^{\mathsf{T}}\right]=\sum_{j=1}^{n}d_{jj}^{2}$ (4)
-where $\operatorname{tr}[.]$ is the matrix trace operator.
-Let $\mathbf{X}_{r}$ be the matrix obtained using the first $r$ terms in the sum of relation (2).
-The Frobenius norm allows to quantify the error of the $r$ rank approximation for $\mathbf{X}$, [15, 18]:
-$\|\mathbf{X}-\mathbf{X}_{r}\|_{\mathsf{F}}^{2}=\sum_{j=r+1}^{n}d_{jj}^{2}$ (5)
-The SVD also allows to solve the linear system of equations which occurs in many engineering problems: $\mathbf{Ax}=\mathbf{b}$, where $\mathbf{A}$ and $\mathbf{b}$ are known matrix and vector of size $m\times n$ and $n$, respectively.
-The entries of the unknown vector $\mathbf{x}$ should be found.
-For a square matrix, $\mathbf{A}$, the Lanczos algorithm, [23], allows to obtain the solution through matrix eigenvalues, see below.
-For a non-square matrix $\mathbf{A}$, the normal equation, $\mathbf{A}^{\mathsf{T}}\mathbf{Ax}=\mathbf{A}^{\mathsf{T}}\mathbf{b}$ which involves a symmetric can be used to search for the vector $\mathbf{x}$.
-For a sparse matrix with large $m$ and $n$ values, the Lanczos algorithm has been extended by Larsen, [25], to bidiagonalization of the non-square matrix leading to perform its SVD.
-For high dimensional data, when the purpose is to get an approximation of the data matrix, it may not be necessary to get all singular values.
-The Lanczos algorithm extension proposed in [1] allows to compute the few first largest or smallest singular values and their associated vectors.
+All singular values and the singular vectors are simultaneously recovered by these approaches.
+1.5 Iterative methods for computing the SVD
+The iterative search methods are used to construct a collection of mutually orthogonal subspaces leading to the singular values and singular vectors.
+From the Rayleigh-Ritz theorem, the minimum and the maximum eigenvalue of a symmetric matrix $\mathbf{A}$ verify ($\mathbf{v}^{\mathsf{T}}\mathbf{v}=1$): $\lambda_{max}(\mathbf{A})\geq\mathbf{v}^{\mathsf{T}}\mathbf{A}\mathbf{v}\geq\lambda_{min}(\mathbf{A})$.
+For a non null vector $\mathbf{v}$, the Rayleigh quotient is: $r(\mathbf{v})=\mathbf{v}^{\mathsf{T}}\mathbf{A}\mathbf{v}/\mathbf{v}^{\mathsf{T}}\mathbf{v}$.
+A gradient search method based on the Rayleigh quotient has been used in [28] to get the eigenvalues and the eigenvectors of the matrix $\mathbf{A}$.
+The Krylov $k$ columns (subspaces) matrix associated with the symmetric matrix $\mathbf{A}$ and the vector $\mathbf{v}$ is: $\mathsf{K}(\mathbf{A},\mathbf{v},k)=(\mathbf{v},\ \mathbf{Av},\ \mathbf{A}^{2}\mathbf{v},\ \ldots,\ \mathbf{A}^{k-1}\mathbf{v})$.
+A next subspace, a column, in this matrix is obtained from the previous by using only a matrix-vector product.
+The Lanczos iterative methods consist of generating Krylov subspaces to get the eigenvalues and the eigenvectors.
+The Arnoldi method uses minimum iterations [1, 33].
+The SVD is also used to solve the linear system of equations which occurs in many engineering problems: $\mathbf{Ax}=\mathbf{b}$, where $\mathbf{A}$ and $\mathbf{b}$ are known matrix and vector of size $m\times n$ and $n$, respectively, and $\mathbf{x}$ is the unknown vector.
+For a square matrix, $\mathbf{A}$, the Lanczos algorithm, [36], allows to get a solution through the EVD.
+For a non-square matrix $\mathbf{A}$, the normal equation, $\mathbf{A}^{\mathsf{T}}\mathbf{Ax}=\mathbf{A}^{\mathsf{T}}\mathbf{b}$, involves a symmetric matrix which can be used to search for $\mathbf{x}$.
+For a sparse matrix with large size, the Lanczos algorithm has been extended by Larsen, [38], to bidiagonalization of the non-square matrix leading to perform its SVD.
+For high dimensional data, when the purpose is to get a low-rank data matrix approximation, it may not be necessary to get all singular values.
+The Lanczos algorithm extension proposed in [2] allows to compute the few first largest or smallest singular values and the singular vectors.
+In the Davidson type iterative methods, polynomial filters are used to get the SVD subspaces, [57, 65].
 1.6 Probabilistic SVD computation
-Performing complete SVD can be computationally costly, especially for high dimensional data.
-When the first important singular values are required, the randomized methods may be used.
-The computation of the SVD for massive data has been considered in [27, 17, 16, 38], where, only the needed $r$ singular values and their associated vectors are computed using a probabilistic (randomized) approach.
-The search consists of finding column orthonormal matrix $\mathbf{Q}$ of size $m\times r$ such that:
+The computation of the SVD for massive data has been considered in [42, 26, 25, 59], where, only the $r$ first largest singular values and singular vectors are computed using a probabilistic (randomized) approach.
+The search consists of finding column orthogonal matrix $\mathbf{Q}$ of size $m\times r$ such that:
 $\|\mathbf{X}-\mathbf{QQ}^{\mathsf{T}}\mathbf{X}\|\leq\epsilon$ (7)
-[…]
+where $\|.\|$ is the $\ell_{2}$-norm, $r$ and $\epsilon$ are settings for the rank and the tolerance of the approximation.
+In the randomized approach, the search for the $r$ first singular values and singular vectors are performed through a random matrix $\boldsymbol{\Omega}$ of size, $n\times(r+\ell)$, where $\ell$ is an integer chosen such that $r+\ell<n$.
+The entries of this matrix are typically from the standard normal distribution: $\boldsymbol{\Omega}=(\omega_{ij})$, $\omega_{ij}\sim\mathcal{N}(0,1)$, $\forall i,j$.
+$\boldsymbol{\Omega}$ is used to define an intermediate matrix, $\mathbf{Y}\triangleq\mathbf{X}\boldsymbol{\Omega}$, which has few columns compared to $\mathbf{X}$.
+A QR decomposition, [24], of the matrix $\mathbf{Y}$ is next performed: $\mathbf{X}\boldsymbol{\Omega}=\mathbf{QR}$, where the columns of the matrix $\mathbf{Q}$ are orthogonal.
+Setting $\mathbf{Z}\triangleq\mathbf{Q}^{\mathsf{T}}\mathbf{X}$ allows to have $\mathbf{QZ}=\mathbf{QQ}^{\mathsf{T}}\mathbf{X}$ or $\mathbf{QZ}=\mathbf{X}$, thank to the column orthogonality of $\mathbf{Q}$.
+Thus, the SVD of the matrix $\mathbf{Z}$ is performed and the result $\mathbf{QZ}$ is returned.
+Indeed, $\mathbf{QZ}=\mathbf{Q}\tilde{\mathbf{U}}\mathbf{DW}^{\mathsf{T}}=\mathbf{UDW}^{\mathsf{T}}$.
+To have rapid singular values decay, the intermediate matrix is computed as: $\mathbf{Y}\triangleq(\mathbf{XX}^{\mathsf{T}})^{q}\mathbf{X}\boldsymbol{\Omega}$.
+Indeed, using the relation $\mathbf{X}=\mathbf{UDW}^{\mathsf{T}}$, the column orthogonal property of the matrices $\mathbf{U}$ and $\mathbf{W}$, and for an integer $q\geq 0$, allow to write:
 $(\mathbf{XX}^{\mathsf{T}})^{q}\mathbf{X}=\mathbf{U}\mathbf{D}^{2q+1}\mathbf{W}^{\mathsf{T}}$ (8)
-This relation shows a ($12q+$ 1)-power of the singular values for this matrix $\mathbf{Y}$.
+This relation shows a ($2q+1$)-power of the singular values for the matrix $\mathbf{Y}$.
 The rapid singular values decay leads to a fast convergence algorithm, [26, 25].
-The paper [17] gives historical overview of the randomized computing and its applications in linear algebra problems.
-Using the probabilistic SVD method, the rank $r$ approximation error is slightly greater than the bound given in relation (5).
-1.3 Eigendecomposition
-The eigendecomposition can be performed for any square matrix. Here, only the symmetric matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ case is considered. The eigendecomposition of this matrix is:
-$\mathbf{X}^{\mathsf{T}}\mathbf{X}=\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{\mathsf{T}}$ (8)
-where $\boldsymbol{\Lambda}$ is a diagonal matrix which entries are nonnegative-valued scalars or eigenvalues: $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}\geq 0$, $\mathbf{V}$ is an orthogonal matrix, $\mathbf{V}^{\mathsf{T}}\mathbf{V}=\mathbf{I}_{n}$, which columns are the eigenvectors associated with the eigenvalues.
-Like in relation (2), the eigendecomposition can be expressed as the sum of rank one matrices or spectral decomposition:
-$\mathbf{X}^{\mathsf{T}}\mathbf{X}=\sum_{j=1}^{n}\lambda_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{\mathsf{T}}$ (9)
-where $\mathbf{v}_{j}$ is the $j$-th column vector of the matrix $\mathbf{V}$.
-The above sum can also be truncated to get an lower rank $r<n$ approximation for the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$.
-The column orthogonal property of the matrix $\mathbf{V}$ allows to have:
-$(\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}=\mathbf{V}\boldsymbol{\Lambda}^{q}\mathbf{V}^{\mathsf{T}}$ (10)
-where $q$ is an integer, $q\geq 1$.
-That means, all power $q$ of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ have the same eigenvectors, only the eigenvalues differ by the power $q$ and we also have: $\lambda_{1}^{q}\geq\lambda_{2}^{q}\geq\ldots\geq\lambda_{n}^{q}\geq 0$.
-At least, the first eigenvalue is nonzero.
-Therefore, we have:
-$(a)$ ::: $1\geq(\lambda_{2}/\lambda_{1})\geq(\lambda_{3}/\lambda_{1})\geq\ldots\geq(\lambda_{n}/\lambda_{1})$ (11)
-$(b)$ ::: $1\geq(\lambda_{2}/\lambda_{1})^{q}\geq(\lambda_{3}/\lambda_{1})^{q}\geq\ldots\geq(\lambda_{n}/\lambda_{1})^{q}$ (12)
-$(c)$ ::: $(\lambda_{j}/\lambda_{1})^{q}\leq(\lambda_{j}/\lambda_{1})\mbox{ ; }j=1,2,\ldots,n$ (13)
-In the principal component analysis, see below, the columns of the matrix $\mathbf{V}$ correspond to projection subspaces.
-One can observe that $-\mathbf{V}$ is also solution of the eigendecomposition in relation (8).
-This observation allows to fulfil the Perron-Frobenius theorem, [31, 12], which stipulates that the entries of the main eigenvector of a nonnegative-valued matrix are all positive or zero.
-Nonnegative-valued matrices occur for many real-life data, e.g.
-genes expression and images.
-The main eigenvector is used in ranking applications, [5, 24].
-To compute the eigenvalue with maximum modulus and the associated eigenvector, the iterative power method, [15], is very often used, see algorithm 2 in the Annexes.
-The convergence rate of the power method depends on the modulus of the second eigenvalue compared to the first, [15, 7].
-At the convergence, the eigenvector is used to get the associated eigenvalue: $\mathbf{X}^{\mathsf{T}}\mathbf{Xv}_{j}=\lambda_{j}\mathbf{v}_{j}$.
-Thank to relation (9), the power method can be used to obtain all eigenvectors and their associated eigenvalues.
+A historical overview of the randomized computing and its applications in linear algebra problems is provided in [26].
+Using the probabilistic SVD method, the $r$ rank approximation error is slightly greater than the bound given below in relation (14), [26].
+In addition, the number of columns, $r+\ell$, of the random matrix $\boldsymbol{\Omega}$ should be large if accurate values are expected for the first $r$ singular values.
+The proposed method leads to more accurate $r$ first singular values by using a $r$ columns random matrix.
+More over, the proposed method can be used to compute the complete SVD even for a small size matrix.
+2 Computing the SVD by using a power method
+Thank to relation (5), the power method can be used to get all eigenvalues and eigenvectors, [24].
 Indeed, let us define $\mathbf{S}_{1}=\mathbf{X}^{\mathsf{T}}\mathbf{X}$.
@@
 That is the Jordan SVD algorithm, [58].
-For a rank deficient matrix $\mathbf{X}$, the number of eigenvalues is less than $n$.
-The power method proposed below allows to compute all required eigenvalues together.
-2 Computing singular values using a power method
-Given a matrix $\mathbf{X}$ of size $m\times n$, $m\geq n$, we first show the link between it’s singular value decomposition factors and the eigendecomposition factors of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$.
-Using the expression (1) of $\mathbf{X}$ in the relation (8) allows to write:
-$\mathbf{X}^{\mathsf{T}}\mathbf{X}=(\mathbf{UDW}^{\mathsf{T}})^{\mathsf{T}}\mathbf{UDW}^{\mathsf{T}}=\mathbf{WD}^{2}\mathbf{W}^{\mathsf{T}}$
-Comparing the above expression to relation (8) allows to express the SVD factors using the eigendecomposition factors:
-[…]
-The entries of the diagonal matrix $\boldsymbol{\Lambda}$ are the diagonal entries of the matrix $(\mathbf{XW})^{\mathsf{T}}\mathbf{XW}$.
-Relation (14) shows a one-to-one relationship between the SVD factors of the matrix $\mathbf{X}$ and the eigendecomposition factors of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$.
-From the above relation, the SVD of the matrix $\mathbf{X}$ can be written as:
+Instead of computing the eigenpairs of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$, one by one, the proposed algorithm computed them simultaneously.
+The column orthogonal property of the matrix $\mathbf{V}$ allows to have:
+$(\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}=\mathbf{V}\boldsymbol{\Lambda}^{q}\mathbf{V}^{\mathsf{T}}$ (9)
+where $q$ is an integer, $q\geq 1$.
+That means, all power $q$ of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ have the same eigenvectors, only the eigenvalues differ by the power $q$ and we also have: $\lambda_{1}^{q}\geq\lambda_{2}^{q}\geq\ldots\geq\lambda_{n}^{q}\geq 0$.
+At least, the first eigenvalue is nonzero, that allows to write:
+$(a)$ $:$ $1\geq(\lambda_{2}/\lambda_{1})\geq(\lambda_{3}/\lambda_{1})\geq\ldots\geq(\lambda_{n}/\lambda_{1})$ (10)
+$(b)$ $:$ $1\geq(\lambda_{2}/\lambda_{1})^{q}\geq(\lambda_{3}/\lambda_{1})^{q}\geq\ldots\geq(\lambda_{n}/\lambda_{1})^{q}$ (11)
+$(c)$ $:$ $(\lambda_{j}/\lambda_{1})^{q}\leq(\lambda_{j}/\lambda_{1})\mbox{ ; }j=1,2,\ldots,n$ (12)
+The proposed power method is based on the Frobenius norm of the matrix $\mathbf{X}$:
+$\|\mathbf{X}\|_{\mathsf{F}}^{2}=\sum_{i=1}^{m}\sum_{j=1}^{n}x_{ij}^{2}=\operatorname{tr}\left[\mathbf{X}^{\mathsf{T}}\mathbf{X}\right]=\operatorname{tr}\left[\mathbf{XX}^{\mathsf{T}}\right]=\sum_{j=1}^{n}d_{jj}^{2}$ (13)
+where $\operatorname{tr}[.]$ is the matrix trace operator.
+Let $\mathbf{X}_{r}$ be the matrix obtained using the $r$ first terms in the sum of relation (2).
+The Frobenius norm allows to quantify the error of the $r$-rank $\mathbf{X}$ approximation, [24, 29]:
+$\|\mathbf{X}-\mathbf{X}_{r}\|_{\mathsf{F}}^{2}=\sum_{j=r+1}^{n}d_{jj}^{2}$ (14)
+Relation (6) shows a one-to-one relationship between the SVD factors of $\mathbf{X}$ and the EVD factors of $\mathbf{X}^{\mathsf{T}}\mathbf{X}$.
+From relation (6), the SVD of the matrix $\mathbf{X}$ can be written as:
 $\mathbf{X}=\mathbf{XWW}^{\mathsf{T}}$ (15)
-When $\mathbf{W}\triangleq\mathbf{Q}$, the above relation is equivalent to that is used in relation (6) and is obtained when performing the eigendecomposition of the matrix $\mathbf{XX}^{\mathsf{T}}$.
+When $\mathbf{W}\triangleq\mathbf{Q}$, the above relation is equivalent to that is used in relation (7) and is obtained when performing the EVD of the matrix $\mathbf{XX}^{\mathsf{T}}$.
 Since we assumed that $n\leq m$, working with the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ necessitates less computational load compared to the matrix $\mathbf{XX}^{\mathsf{T}}$.
 To obtain the column orthogonal matrix $\mathbf{W}$ given $\mathbf{X}$, a norm of the difference between the left-hand and right-hand side terms in the above relation can be minimized.
-In the probabilistic SVD approach, a column orthonormal matrix of size $r+\ell$, $\ell>r$, is used and the QR decomposition is performed followed by matrices multiplication to get the SVD result.
-Here, the entries of a random column orthogonal matrix of size $n\times r$ are iteratively modified to fit the factor $\mathbf{W}$ for the SVD result of the matrix $\mathbf{X}$.
-The algorithm used is similar to the power method but all eigenvectors are simultaneously recovered at a time.
-A new parameter is introduced to increase the convergence rate, compared to the basic power method.
-2.1 Search for the projection matrix using the gradient method
+In the probabilistic SVD approach, a column orthogonal matrix of size $\ell, >r\ell>r$, is used and the QR decomposition is performed followed by matrices multiplication to get the SVD.
+Here, the entries of a random column orthogonal matrix of size $n\times r$ are iteratively modified to fit the factor $\mathbf{W}$ of the SVD.
+The algorithm used is similar to the power method but the $r\leq n$ eigenvectors are simultaneously recovered.
+2.1 Search for the projection matrix by using the gradient method
 To obtain the matrix $\mathbf{W}$ in relation (15), the Frobenius norm of the difference between $\mathbf{X}$ and $\mathbf{XWW}^{\mathsf{T}}$ is minimized, subject to the column orthogonality constraint for the matrix $\mathbf{W}$.
-The following constrained optimization problem is solved:
-$\min\mathsf{J}(\mathbf{W})\mbox{, }\mathsf{J}(\mathbf{W})=\frac{1}{2}\left\|\mathbf{X}-\mathbf{XWW}^{\mathsf{T}}\right\|_{\mathsf{F}}^{2}\mbox{ subject to: }\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}$ (16)
-where .$\|.\|_{\mathsf{F}}^{2}$ is the square of the Frobenius norm. The Lagrangian of this optimization problem is:
+$\min\mathsf{J}(\mathbf{W})\mbox{; }\mathsf{J}(\mathbf{W})=\frac{1}{2}\left\|\mathbf{X}-\mathbf{XWW}^{\mathsf{T}}\right\|_{\mathsf{F}}^{2}\mbox{ subject to: }\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}$ (16)
+The Lagrangian of this optimization problem is:
 $\mathcal{L}(\mathbf{W},\Upsilon)=\frac{1}{2}\left\|\mathbf{X}-\mathbf{XWW}^{\mathsf{T}}\right\|_{\mathsf{F}}^{2}-\left\|\Upsilon\left(\mathbf{W}^{\mathsf{T}}\mathbf{W}-\mathbf{I}\right)\right\|_{\mathsf{F}}^{2}$ (17)
 where the entries of the matrix $\Upsilon$ are the Lagrange multipliers associated with the equality constraints.
-To minimize $\mathsf{J}(\mathbf{W})$, the gradient descent method can be used with the column orthogonal constraint for the matrix $\mathbf{W}$.
-For the gradient search optimization method, the solution at iteration $t$ is obtained after updating the previous one: $\mathbf{W}^{(t)}=\mathbf{W}^{(t-1)}-\eta\nabla\mathcal{L}(\mathbf{W},\Upsilon)$, where $\eta$ is a positive step parameter and $\nabla\mathcal{L}(\mathbf{W},\Upsilon)$ is the gradient of the Lagrangian.
-To obtain the matrix $\mathbf{W}$ leading to minimum of the Lagrangian, the derivative of $\mathcal{L}(\mathbf{W},\Upsilon)$ with respect to $\mathbf{W}$ and $\Upsilon$ should be computed.
+To minimize $\mathsf{J}(\mathbf{W})$, the gradient descent method is used, the solution at iteration $t$ is obtained after updating the previous one: $\mathbf{W}^{(t)}=\mathbf{W}^{(t-1)}-\eta\nabla\mathcal{L}(\mathbf{W},\Upsilon)$, where $\eta$ is a positive step parameter and $\nabla\mathcal{L}(\mathbf{W},\Upsilon)$ is the gradient of the Lagrangian.
+The derivative of $\mathcal{L}(\mathbf{W},\Upsilon)$ with respect to $\mathbf{W}$ and $\Upsilon$ are then required.
 Relation (13) allows to write:
@@
 $=$ $\operatorname{tr}\left[\mathbf{X}\mathbf{X}^{\mathsf{T}}-2\mathbf{XWW}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}+\mathbf{XWW}^{\mathsf{T}}\mathbf{WW}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}\right]$
-For any column orthogonal matrix $\mathbf{W}$ of size $n\times r$, $\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{r}$, we then have:
+For any column orthogonal matrix $\mathbf{W}$ of size $n\times r$, $\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{r}$, that allows to write:
 $\left\|\mathbf{X}-\mathbf{XWW}^{\mathsf{T}}\right\|_{\mathsf{F}}^{2}$ $=$ $\operatorname{tr}\left[\mathbf{X}\mathbf{X}^{\mathsf{T}}-\mathbf{XWW}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}\right]$ (18)
-$\left\|\mathbf{W}^{\mathsf{T}}\mathbf{W}-\mathbf{I}\right\|_{\mathsf{F}}^{2}$ $=$ 00 0 (19)
-For an $n$-order orthogonal matrix $\mathbf{W}$, we also have $\mathbf{W}\mathbf{W}^{\mathsf{T}}=\mathbf{I}_{n}$.
+$\left\|\mathbf{W}^{\mathsf{T}}\mathbf{W}-\mathbf{I}\right\|_{\mathsf{F}}^{2}$ $=$ $0$ (19)
+For any $n$-order orthogonal matrix $\mathbf{W}$, we have $\mathbf{W}\mathbf{W}^{\mathsf{T}}=\mathbf{I}_{n}$.
 However, all $n$-order orthogonal matrices do not allow to verify relation (4) for a given matrix $\mathbf{X}$.
@@
 The matrix $\mathbf{XX}^{\mathsf{T}}-\mathbf{XWW}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}$ is symmetric and has nonnegative-valued eigenvalues.
-The trace of this matrix is obtained as the sum of its eigenvalues.
-For an $n$-order column orthogonal matrix $\mathbf{W}$, relation (18) shows that the global minimum of the objective function $\mathsf{J}(\mathbf{W})$ is zero.
-For an $r$-column orthogonal matrix $\mathbf{W}$, $r<n$, we have:
+The trace of this matrix is the sum of its eigenvalues.
+For a $n$-order column orthogonal matrix $\mathbf{W}$, relation (18) shows that the global minimum of the objective function $\mathsf{J}(\mathbf{W})$ is zero.
+For a $r$-column orthogonal matrix $\mathbf{W}$, $r<n$, we have:
 $tr\left[\mathbf{XX}^{\mathsf{T}}-\mathbf{XWW}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}\right]=tr\left[\mathbf{XX}^{\mathsf{T}}\right]-tr\left[\mathbf{XW}(\mathbf{XW})^{\mathsf{T}}\right]=\sum_{j=r+1}^{n}d_{jj}$
 where $\mathbf{XW}(\mathbf{XW})^{\mathsf{T}}$ is the diagonal matrix which entries are the $r$ first singular values, see the text under relation (6).
-The relation (118) in Petersen & Petersen [32, page 13], can be written as:
+The relation (118) in [51, page 13], can be written as:
 $\frac{\partial}{\partial\mathbf{W}}\operatorname{tr}\left[\mathbf{AWBW}^{\mathsf{T}}\mathbf{C}\right]=\mathbf{A}^{\mathsf{T}}\mathbf{C}^{\mathsf{T}}\mathbf{WB}^{\mathsf{T}}+\mathbf{CAWB}$
@@
 $\frac{\partial\mathsf{J}(\mathbf{W})}{\partial\mathbf{W}}=-\mathbf{X}^{\mathsf{T}}\mathbf{XW}$
-The matrix $\mathbf{W}$ in the above relation is that available in the previous iteration: $\mathbf{W}^{(t-1)}$.
+The matrix $\mathbf{W}$ in the above relation is that available in the previous iteration, $\mathbf{W}^{(t-1)}$.
 Then, the gradient updating equation is: $(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})\mathbf{W}^{(t-1)}$.
-Therefore, at iteration $t=1,2,\ldots$ and given column orthogonal matrix $\mathbf{W}^{(t-1)}$, the gradient descent optimization search method consists of two alternate steps to obtain the update:
+Therefore, at iteration $t=1,2,\ldots$ and given column orthogonal matrix $\mathbf{W}^{(t-1)}$, two alternate steps are used to get the update $\mathbf{W}^{(t)}$:
 $\tilde{\mathbf{W}}$ $=$ $\mathbf{G}\mathbf{W}^{(t-1)}$ (20)
-$\mathbf{W}^{(t)}$ $=$ $Gram-Schmidt(\tilde{\mathbf{W}})$ (21)
-where $\mathbf{G}\triangleq(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$, $\eta$ is a positive step parameter and $q\geq 1$ is an integer.
+[…]
+where $\mathbf{G}\triangleq\mathbf{I}_{n}+\eta(\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$, $\eta$ is a positive step parameter and $q\geq 1$ is an integer.
 The settings for $\eta$ and $q$ are discussed later.
 The method to use for obtaining the column orthogonal matrix $\mathbf{W}$ is summarized in algorithm 1.
-Data: matrix $\mathbf{X}$ of size $m\times n$, ($m\geq n$), settings $\eta$, $q$ and $\epsilon$;
-Result: matrix $\mathbf{W}$ with orthogonal columns;
-1 Initialization: select randomly entries of a matrix $\mathbf{W}$ from a normal or an uniform distribution and perform orthogonalization $\mathbf{W}^{(0)}=$ Gram-Schmidt($\mathbf{W}$), compute $\mathbf{G}=(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$;
-2 for $t\leftarrow 1,2,\ldots$ do
+Data: matrix $\mathbf{X}$ of size $m\times n$, settings $\eta$, $q$, $\epsilon$ and maximum number of iterations (itmaxitmax);
+Result: column orthogonal matrix $\mathbf{W}$;
+1 Initialization: select randomly entries of a matrix $\mathbf{W}$ from a normal or an uniform distribution and perform orthogonalization $\mathbf{W}^{(0)}=$ Gram-Schmidt($\mathbf{W}$), compute $\mathbf{G}=\mathbf{I}_{n}+\eta(\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$;
+2 for $t\leftarrow 1,2,\ldots,itmax$ do
 3   compute $\tilde{\mathbf{W}}$ using relation (20);
@@
 10 end for
-Algorithm 1 Gradient search for columns orthonormal matrix $\mathbf{W}$
-This algorithm is similar to the power method, see below.
-2.1.1 Factors for the eigendecomposition of $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and for the SVD of $\mathbf{X}$
-The algorithm 1 leads to a column orthogonal matrix $\mathbf{W}$ that fits the matrix $\mathbf{X}$.
-Using relation (8), the eigenvalues of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ are the entries of the diagonal matrix $\boldsymbol{\Lambda}$.
-These eigenvalues are the diagonal entries of the matrix $(\mathbf{XW})^{\mathsf{T}}\mathbf{XW}$ which is symmetric.
-The diagonal matrix of the SVD is: $\mathbf{D}=\boldsymbol{\Lambda}^{1/2}$.
-Finally the matrix $\mathbf{U}$ of the SVD is: $\mathbf{U}=\mathbf{XW}\boldsymbol{\Lambda}^{-1/2}$.
-2.1.2 Relation with the power method
-Algorithm 1 is in fact the power method allowing to obtain all eigenvalues together and not one by one as done using the classical power method algorithm 2, see Annexes.
-The matrix $\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)$ is used in this algorithm instead of the matrix $\mathbf{A}\triangleq\mathbf{X}^{\mathsf{T}}\mathbf{X}$.
-These two matrices have the same eigenvectors.
-Indeed, the eigenvector $\mathbf{z}$ associated with the eigenvalue $\gamma$ of the matrix $\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)$ allows to write:
-$(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})\mathbf{z}=\gamma\mathbf{z}\Rightarrow\mathbf{X}^{\mathsf{T}}\mathbf{Xz}=\left(\frac{\gamma-1}{\eta}\right)\mathbf{z}$
-The above relation means the matrices $\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)$ and $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ have the same eigenvectors which are associated with related eigenvalues, $j=1,2,\ldots,n$:
+Algorithm 1 Gradient search for column orthogonal matrix $\mathbf{W}$
+The columns of the result matrix $\mathbf{W}$ are the eigenvectors matrix of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and the right singular vectors matrix of $\mathbf{X}$.
+Using relation (4), the eigenvalues of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ are the entries of the diagonal matrix $\boldsymbol{\Lambda}$ or the diagonal entries of the symmetric matrix $(\mathbf{XW})^{\mathsf{T}}\mathbf{XW}$.
+Relation (6) allows to get the SVD factor matrices $\mathbf{D}$ and $\mathbf{U}$.
+2.2 Relation with the power method
+Algorithm 1 is a power method allowing to get the $r\leq n$ eigenvectors simultaneously and not one by one like by the classical power method in algorithm 2, see Annexes.
+The matrix $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$ is used in algorithm 1 instead of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$.
+Indeed, the eigenvector $\mathbf{z}$ associated with the eigenvalue $\gamma$ of the matrix $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$ allows to write:
+$\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)\mathbf{z}=\gamma\mathbf{z}\Rightarrow\mathbf{X}^{\mathsf{T}}\mathbf{X}\mathbf{z}=\left(\frac{\gamma-1}{\eta}\right)\mathbf{z}$
+The above relation means the matrices $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ have the same eigenvectors which are associated with related eigenvalues, $j=1,2,\ldots,n$:
 $\lambda_{j}$ $=$ $(\gamma_{j}-1)/\eta$ (22)
@@
 where $\lambda_{j}$ and $\gamma_{j}$ are the eigenvalues of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$, respectively.
-Relation (23) shows that the smallest eigenvalue of the matrix $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$ is one, meaning this matrix is always of full rank even for rank deficient matrix $\mathbf{X}$.
-Safety computation of the norm $\|\mathbf{W}^{(t)}-\mathbf{W}^{(t-1)}\|_{\mathsf{F}}^{2}$, step 5 of alogithm 1, at any iteration requires to have the same number of columns for the matrix $\mathbf{W}$.
+Relation (23) shows that the smallest eigenvalue of the matrix $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$ is one, meaning this matrix is always of full rank even for a rank deficient matrix $\mathbf{X}$.
+A safe computation of the norm $\|\mathbf{W}^{(t)}-\mathbf{W}^{(t-1)}\|_{\mathsf{F}}^{2}$, step 5 of alogithm 1, at any iteration requires to have the same number of columns for the matrix $\mathbf{W}$.
 When the matrix $\mathbf{X}$ is not of full rank, there exit a matrix $\mathbf{W}$ of size $n\times r$, $r<n$, allowing to have an exact reconstruction of the $\mathbf{X}$.
-Therefore, using alone the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ in algorithm 1 may lead a problem for computing the above norm if we are not careful.
-2.1.3 Convergence analysis and choice of the settings $\eta$ and $q$
+Therefore, using alone the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ in algorithm 1 may lead to a problem for computing the above norm if we are not careful.
+2.3 Convergence analysis and choice of the settings $\eta$ and $q$
 From the convex property of the objective function, the algorithm 1 is convergent.
-The rate of convergence depends especially on the behavior of the singular values of the matrix $\mathbf{X}$ and on the precision ($\epsilon$) required.
+The rate of convergence depends on the behavior of the singular values of the matrix $\mathbf{X}$, the precision $\epsilon$ and the step parameter $\eta$.
 The rate of convergence for the probabilistic SVD is analyzed in [26, 25] where faster convergence is observed for matrix which singular values decay quickly.
@@
 This result corresponds to the modulus conditions for the power method.
-Relation (13) shows a possibility to modify the singular values decay using power $q$ of the matrix in the eigendecomposition.
-From relation (23), a small value (near zero) for the parameter $\eta$ leads to have all eigenvalues for matrix $\mathbf{G}$ close to one.
+Relation (12) shows a possibility to modify the singular values decay by using a power $q$ of the matrix in the EVD.
+From relation (23), a small $\eta$ value (near zero) may lead to near one eigenvalues for the matrix $\mathbf{G}$.
 On the other hand, a value greater to $1$ for the parameter $\eta$ leads to increasing the eigenvalues of the matrix $\mathbf{G}$.
-Therefore, the choices $q>1$ for the power parameter and $\eta\geq 1$ for the step parameter, likely lead to rapid singular values decay for the matrix $\mathbf{G}$.
-The computation of the matrix $\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)^{q}$ may lead to numerical problem when $\mathbf{X}$ has some entries with very small or high values.
-This square matrix of order $n$ is computed once before starting the iterative search.
+Therefore, the choices $q>1$ and $\eta\geq 1$, likely lead to rapid singular values decay for the matrix $\mathbf{G}$.
+However, when the matrix $\mathbf{X}$ has large eigenvalues, small $\eta$ value may be good.
+The matrix $\mathbf{G}=\mathbf{I}_{n}+\eta(\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$ which is computed once before starting the iterative search, may lead to numerical problem when $\mathbf{X}$ has some very small or high entries.
 At each iteration, a matrix multiplication and a Gram-Schmidt (GS) orthogonalization are performed.
-The modified GS orthogonalization method is summarized in algorithm 3, in the Annexes.
-The matrices multiplication and the GS orthogonalization are the time-consuming part of the proposed method which is implemented in the psvd R package of the R environment, [35] (https://cran.r-project.org/web/packages/psvd/index.html).
-A C code is used to speed-up the matrices multiplication and the GS orthogonalization calculator.
-2.1.4 Small size matrice illustrative examples
-To illustrate the choice of settings, three academic example matrices, relation (24), and the Fisher’s iris data, [11], were used.
-The iris dataset corresponds to measurements of the length and the width of the sepal and the petal for $3$ iris species: setosa, versicolor and virginica.
+The modified GS orthogonalization method is summarized in algorithm 3, see Annexes.
+The matrices multiplication and the GS orthogonalization are the heavy computing time part of the proposed method which is implemented in the psvd R, [54], package (https://cran.r-project.org/web/packages/psvd/index.html).
+A C code is used to speed-up the matrices multiplication and the GS orthogonalization.
+3 Experimental results
+The implementation of algorithm 1 in the R package psvd has the following inputs: $X$ the matrix, $r$ (the first singular values, default is 2), $\eta$ (the gradient search positive scalar, default is 1), $q$ (power integer, default is 2), itmax (maximum iterations number, default is 200), errerr (convergence error default is 1e-8) and mySeed (a seed used for generating an initial solution, default is 50).
+The matrix $X$ is mandatory and $r$ is a user supply parameter.
+The choice of the parameters $\eta$ and $q$ are discussed in subsection 2.3.
+A large number of iterations number itmax may lead to accurate singular values, especially a large size data matrix where the convergence can be slow or for a challenging matrix example.
+The parameter err is used for convergence, see step 6 of algorithm 1.
+The parameter mySeed allows to have reproducible results for different runs on the same matrix.
+To show the effectiveness of the proposed SVD method and to illustrate the choice of the settings, some numerical examples are used.
+3.1 Small size matrix examples
+Three academic matrix examples, relations (24)-(25), and the Fisher’s iris data, [18], were used.
+The iris dataset corresponds to the length and the width measurements of the sepal and the petal for $3$ iris species: setosa, versicolor and virginica.
 $50$ measurements were observed for each species leading to $m=150$ and $n=4$.
-$X_{a}=\left(\begin{array}[]{ccc}1&1&1\\ 0&2&1\\ 1&0&1\end{array}\right)\mbox{ ; }X_{b}=\left(\begin{array}[]{cccc}3&1&9&2\\ 10&4,&8&6\\ 7&6&12&1\\ 11&2&5&9\\ 1&1&1&0\end{array}\right)\mbox{ ; }X_{c}=\left(\begin{array}[]{ccccc}22&10&2&3&7\\ 14&7&10&0&8\\ -1&13&-1&-11&3\\ -3&-2&13&-2&4\\ 9&8&1&-2&4\\ 9&1&-7&5&-1\\ 2&-6&6&5&1\\ 4&5&0&-2&2\end{array}\right)$ (24)
-The singular values associates with these examples are shown in Table 1. The second and the third eigenvalues for the matrix $\mathbf{X}_{c}$ are very close and the last two eigenvalues are zero.
-Table 1: Singular values
+$X_{a}=\left(\begin{array}[]{ccc}1&1&1\\ 0&2&1\\ 1&0&1\end{array}\right)\mbox{ ; }X_{b}=\left(\begin{array}[]{cccc}3&1&9&2\\ 10&4,&8&6\\ 7&6&12&1\\ 11&2&5&9\\ 1&1&1&0\end{array}\right)$ (24)
+$X_{c}=\left(\begin{array}[]{ccccc}22&10&2&3&7\\ 14&7&10&0&8\\ -1&13&-1&-11&3\\ -3&-2&13&-2&4\\ 9&8&1&-2&4\\ 9&1&-7&5&-1\\ 2&-6&6&5&1\\ 4&5&0&-2&2\end{array}\right)$ (25)
+The singular values associates with these four examples are shown in Table 1.
+The second and the third eigenvalues for the matrix $\mathbf{X}_{c}$ are very close and the last two eigenvalues are zero.
+The matrix $\mathbf{X}_{c}$ is a challenging example for many SVD computation algorithms.
+Table 1: Singular values for the matrices in relations (24) - (25) and for the iris dataset
 # $X_{a}$ $X_{b}$ $X_{c}$ Iris
 1 2.80193774 26.02508484 35.32704347 95.95991387
-2 1.44504187 9.31733797 20 17.76103366
-3 0.24697960 3.29881377 19.59591794 3.46093093
-4 - 0 0 1.88482630
-5 - - 0 -
-Table 2: Convergence iterations
-$q=1$
-$\eta$ 0.01 0.1 1 10 100
-$X_{a}$ 500 81 20 16 16
-$X_{b}$ 156 25 11 11 11
-$X_{c}$ 420 345 342 341 341
-Iris 231 42 21 18 18
+2 1.44504187 9.317337969 20.00000000 17.76103366
+3 0.24697960 3.298813773 19.59591794 3.460930930
+4 - 0.000000000 0.000000000 1.884826306
+5 - - 0.000000000 -
+Table 2: Iterations number for small size matrices examples, the maximum number is set to 200
 $q=2$
-$\eta$ 0.01 0.1 1 10 100
-$X_{a}$ 328 43 11 9 9
-$X_{b}$ 82 14 6 6 6
-$X_{c}$ 221 184 180 179 184
-iris 120 22 11 10 10
-For fixed power parameter $q=1,2$, the step parameter $\eta$ was varied.
-The number of iterations required for $8$ digits convergence to the values in Table 1 are shown in Table 2.
+$\eta$ 0.01 0.1 1 2 5 10 100
+$X_{a}$ 201 43 11 9 9 9 9
+$X_{b}$ 24 9 6 6 6 6 6
+$X_{c}$ 179 179 179 179 179 179 179
+Iris 27 12 10 10 10 10 10
+$q=3$
+$\eta$ 0.01 0.1 1 2 5 10 100
+$X_{a}$ 160 26 9 7 6 6 6
+$X_{b}$ 8 5 5 5 5 5 5
+$X_{c}$ 123 123 125 124 133 145 201
+iris 10 8 8 7 7 8 7
+$q=4$
+$\eta$ 0.01 0.1 1 2 5 10 100
+$X_{a}$ 85 16 7 6 5 7 7
+$X_{b}$ 5 4 4 4 4 4 4
+$X_{c}$ 140 178 201 201 201 201 201
+iris 39 201 201 201 201 201 106
+For a power parameter $q\in\{2,3,4\}$, the step parameter $\eta$ was varied.
+The number of iterations required for convergence, $8$ digits after the decimal point, are shown in Table 2 where $201$ means no convergence after $200$ iterations.
+For the examples $X_{a}$ and $X_{b}$, rapid convergence is observed by using $\eta\geq 1$ and $q\in\{2,3,4\}$.
+The settings $\eta\leq 0.1$ and $q=3$ seem appropriate for the example $X_{c}$ while $q=3$ and $\eta\geq 1$ are better for the iris dataset.
 Using $q=3$ or $q=4$ necessitates the same number of operations in the initialization step.
-For some test examples not shown, using $q=4$ do not show a significant reduce of the number of iterations than using $q=2$.
-In addition, using $q=4$ may lead to huge entries in the matrix $(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$ for some examples.
-Therefore, the choice $q=2$ is a good compromise.
-2.1.5 High-dimensional example: trachea data
+Using $q=4$ may lead to huge entries in the matrix $\mathbf{I}_{n}+\eta(\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$ for some data examples.
+Therefore, the choics $q=2,3$ are good compromise.
+The choice $q=3$ may be appropriate for large size dataset and it may be useful to scale the matrix entries before computing the SVD.
+The scaling of a nonnegative-valued matrix can consist of dividing each entry by the maximum value to have a matrix which entries vary in the interval $[0,1]$.
+The singular values are then divided by the same scalar.
+3.2 High-dimensional example: trachea data
 The trachea dataset corresponds to gene expression profiling using the sequencing technology.
 The data was downloaded from the Gene Expression Omnibus, [5], using the access number GSE103354, file: GSE103354_Trachea_fullLength_TPM.txt.gz.
-This dataset consists of full-length scRNA-seq for $301$ trachea segments of C57BL/6 wild-type mice, [29].
+This dataset is the full-length single cell RNA sequencing for $301$ trachea segments of C57BL/6 wild-type mice, [46].
 A total of $23,420$ genes (transcripts) are reported.
 There are $4,836$ transcripts which have zero count values for the $301$ cells.
-These transcripts were removed and the remaining dataset has $18,584$ rows.
-Figure 1: Log scale eigenvalues of the trachea dataset
-Figure 1 shows the behavior of the $\log$ scale singular values.
-For this example, the first $87$ singular values capture $90\%$ of information and the first $159$ singular values capture $95\%$ of information.
-One can observe a fast decay for the first $20$ singular values and a slow decay for the others.
+These transcripts (rows) were removed and the remaining sparse dataset has $18,584$ rows, $58\%$ of zero-valued entries, few entries ($0.52\%$) of value greater than $1,000$ and a maximum entry value $609,926.3$.
+Figure 1: Log scale singular values of the trachea dataset
+Figure 1 shows a $\log$ scale singular values behavior for the trachea dataset.
+There is a fast decay for the $20$ first singular values and a slow decay for the others.
+For this example, the $87$ and $159$ first largest singular values capture $90\%$ and $95\%$ of information, respectively.
 Five methods available in the literature were used to show the performance of the proposed method.
-The calculations were performed on the same laptop computer equipped with i5-1135G7 processor, 8 GB of RAM, under Microsoft Windows 11 family and R environment version 4.4.1, [35].
+The calculations were performed on the same laptop computer equipped with i5-1135G7 processor, 8 GB of RAM, under Microsoft Windows 11 family and R environment version 4.5.1, [54].
 The psvd R package of the proposed method is available from the comprehensive R archive network, http://cran.r-project.org/web/packages/psvd/.
-The classical SVD method is available through the R build-in svd() function which computes all singular values.
+The classical SVD method is available through the R build-in svd() function.
 The Lanczos algorithm based method for computing all singular values is available in the svd R package, [34].
 A C++ library for large scale eigenvalue problem is available in the Spectra tool, https://github.com/yixuan/spectra.
-The RSpectra R package allows to use this library for computing the first $r$ singular values, [34].
-The irlba R package, [2], also allows to compute the first $r$ singular values.
+The RSpectra R package allows to use this library for computing the $r$ first singular values, [53].
+The Lanczos bidiagonalization based irlba R package, [3], also allows to compute the $r$ first singular values.
 The probabilistic SVD method is implemented in the rsvd R package, [15].
-When possible, the largest $20$, $50$, $100$ or $150$ singular values were computed.
-Each method was run $10$ times using default settings and the average results were reported.
-For the proposed method, the following settings were used: $q=2$, tolerance ϵ=italic-$\epsilon=$1e-8 and $\eta=10$.
-The classical svd() function was used as reference and three parameters were examined: the runtime, percentage rate of reconstruction and the means squared error (MSE).
-Table 3: Comparison results
-Method runtime (s) reconst. Rate % MSE
-$r=20$ svd 3.77 80.83 -
-propack.svd 4.58 80.83 1.39e-8
-RSpectra 0.53 80.83 1.53e-8
-irlba 0.42 80.83 0.00160
-rsvd 1.04 80.83 320.106
-psvd 1.66 80.83 2.35e-8
-$r=50$ svd 3.77 86.21 -
-propack.svd 4.58 86.21 1.39e-8
-RSpectra 1.29 86.21 1.56e-8
-irlba 1.36 86.21 0.00554
-rsvd 2.42 86.15 1740.22
-psvd 2.61 86.21 7.77496
-$r=100$ svd 3.77 91.11 -
-propack.svd 4.58 91.11 1.39e-8
-RSpectra 2.12 91.11 0.69e-8
-irlba 5.42 91.11 0.00989
-rsvd 4.92 91.01 2162.08
-psvd 4.50 91.11 38.2471
-$r=150$ svd 3.77 94.47 -
-propack.svd 4.58 94.47 1.39e-8
-RSpectra 3.18 94.47 1.51e-8
-irlba 14.53 94.47 0.01732
-rsvd 7.54 94.38 1688.06
-psvd 6.93 94.47 46.3998
+The $20$, $50$, $100$ or $150$ first largest right singular vectors were computed for $5$ methods, see Table 3.
+The svd, propack.svd, RSpectra, and the psvd methods are almost deterministic and each leads to the same singular values for different runs on the same matrix.
+A small error, $\sim$1e-3, is observed for the irlba method singular values for different runs on the same matrix, this error is larger, ¿10, for the rsvd method.
+These later two methods were run $10$ times, the average results for the irlba method and the best run (minimum mean squares error) results for the psvd method are reported.
+Each method is run with the default settings.
+For the proposed method, the following settings were used: qq=2, tolerance $\epsilon$=1e-8, $\eta$=1 and itmax=200.
+The classical svd() function was used as reference and three parameters were examined: the runtime, percentage rate of reconstruction and the mean squared error (MSE).
+The MSE value is obtained by computing the differences between the $r$ first singular values of a given method and those of the classical SVD method.
+Then, the mean of the squares of the differences is computed and shown.
 The results obtained are summarized in Table 3 where the proposed method appears in bold.
-Runtime.
-All singular values are computed togheter using the classical svd and the propack.svd methods.
-The former method is slightly faster than the later.
-When the first $20$ are required, the RSpectra, irlba, rsvd and psvd methods use less runtime compare to the svd and propack.svd methods.
+Table 3: Comparison of the runtime, the reconstruction rate and the MSE for six SVD methods
+$r$ Method runtime (s) reconst.rate % MSE
+2020 svd 4.12 80.83 -
+propack.svd 0.95 80.83 <<1.0e-10
+RSpectra 0.51 80.83 <<1.0e-10
+irlba 0.41 80.83 2.76e-8
+rsvd 0.98 80.83 889.4
+psvd 1.61 80.83 <<1.0e-10
+$50$ svd 4.12 86.21 -
+propack.svd 2.81 86.21 <<1.0e-10
+RSpectra 1.27 86.21 <<1.0e-10
+irlba 1.34 86.21 3.43e-7
+rsvd 2.14 86.16 33,427.6
+psvd 2.51 86.21 1.21
+$100$ svd 4.12 91.11 -
+propack.svd 3.75 91.11 <<1.0e-10
+RSpectra 2.08 91.11 <<1.0e-10
+irlba 5.42 91.11 4.03e-7
+rsvd 4.51 91.01 36,076.5
+psvd 4.37 91.11 14.63
+$150$ svd 4.12 94.47 -
+propack.svd 3.94 94.47 <<1.0e-10
+RSpectra 3.12 94.47 <<1.0e-10
+irlba 14.22 94.47 5.3e-3
+rsvd 7.47 94.38 16,203.4
+psvd 6.52 94.47 14.35
+Runtime
+All singular values are simultaneously computed using the classical svd method.
+When the first $20$ are required, the RSpectra, irlba, rsvd, propack.svd and psvd methods use less runtime compare to the svd method.
 For this example, when $r=150$, the RSpectra is the fastest followed by classical svd, propack.svd, psvd, rsvd and irlba.
-The runtime of this later method grows quickly when the number of the required singular values increases.
-Reconstruction rate.
+The runtime of this later method grows quickly when the required number of the singular values increases.
+Reconstruction rate
 Relation (3) was used to compute the reconstruction rate.
-The denominator is the same for all method and its value is obtained using the classical svd method results.
-The numerator is computed using the results of each method.
-The reconstruction rate is relatively the same for all methods.
+The denominator is the same for all methods and its value is obtained using the classical svd method result.
+The numerator is computed using the result of each method.
+The reconstruction rate is almost the same for all methods.
 A very small difference appears for the rsvd method when $r\geq 50$.
-Mean squared error.
-The differences between the singular values of each method and those obtained using the classical method where computed and used to calculate a mean squared error (MSE).
+Mean squared error
+The differences between the singular values of each method and those obtained using the classical method where computed and used to calculate the MSE.
 The propack.svd and the RSpectra methods have quasi similar reconstruction rate as the classical svd method.
+The first $20$ singular values obtained by all methods are relatively accurate.
 The irlba method has near zero MSE value.
-The methods psvd and rsvd have more than zero MSE values.
-The psvd method MSE can be improved by using small tolerance $\epsilon$ value.
-The rsvd method has relatively higher MSE value and can lead to a significant gap between singular values, see Table 4 showing the last $10$ singular values for $r=100$.
-From the results of the rsvd method, we observed that the first $50$ estimates obtained using $r=100$ are more accurate than those obtained using $r=50$.
-The first $20$ singular values obtained by all methods are relatively accurate.
-Table 4: $91$-th to $100$-th singular values
-i svd psvd rsvd
-91 10847.51 10846.65 10390.96
-92 10804.19 10805.04 10331.54
-93 10638.48 10638.33 10260.26
-94 10600.40 10600.56 10040.75
-95 10413.53 10413.28 9963.21
-96 10376.98 10377.18 9920.82
-97 10332.29 10332.33 9820.13
-98 10289.21 10289.21 9691.56
-99 10195.62 10195.62 9639.37
-100 10002.21 10002.21 9432.25
-The irlba and the rsvd methods cannot be used for small size matrices. Therefore, only RSpectra and the proposed psvd methods can be used to compute the $r$ largest singular values for any matrix.
-3 Applications
-In the above presentation, the eigendecomposition and the SVD consist of linear algebra manipulations of the matrix $\mathbf{X}$ entries.
-Truncating of the sum in (2) to $r<n$ leads to a $r$ rank approximation of the matrix $\mathbf{X}$.
-A direct application of the $r$ largest eigenvectors associated with the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ is the principal component analysis.
+The psvd and rsvd methods have more than zero MSE values.
+The rsvd method has relatively higher MSE value and can lead to a significant gap between singular values, see Table 4 showing $10$ singular values when $r=100$ (a) or when $r=150$ (b).
+Table 4: $91$-th to $100$-th singular values obtained using the classical svd, the proposed psvd and the rsvd methods: (a) $r=100$, (b) $r=150$
+i svd psvd(a) rsvd(a) rsvd(b)
+91 10,847.51 10,846.65 10,483.18 10,808.09
+92 10,804.19 10,805.04 10,426.29 10,754.47
+93 10,638.48 10,638.33 10,298.68 10,603.37
+94 10,600.40 10,600.56 10,194.08 10,540.89
+95 10,413.53 10,413.28 10,075.35 10,371.87
+96 10,376.98 10,377.18 9,942.93 10,329.69
+97 10,332.29 10,332.33 9,805.83 10,264.57
+98 10,289.21 10,289.21 9,619.66 10,225.44
+99 10,195.62 10,195.62 9,524.24 10,137.75
+100 10,002.21 10,002.21 9,474.41 9,916.85
+MSE - 0.161 231,365 3,206
+Because of the random property of the rsvd method, the values shown in Table 4 for this method may be different for different runs while those for the classical method and the proposed methods remain unchanged.
+The irlba and the rsvd methods should be used for only large size matrices.
+These $2$ methods and the RSpectra method are the fastest for computing a small number of first largest singular vectors.
+Only the propack.svd, the RSpectra and the proposed psvd methods can be used to compute the $r$ largest singular vectors for any matrix.
+For all methods in the comparison, only the $k$ first singular vectors of the proposed method are exactly a neural network autoencoder weight matrix columns.
+Despite its relatively long runtime, the proposed method may be useful to have a direct link between data singular vectors and the columns of a neural network weight matrix.
+4 Application results
+Truncating of the sum in relation (2) to $r<n$ leads to a $r$ rank matrix $\mathbf{X}$ approximation.
+The $r$ first largest eigenvectors of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ are used in the principal component analysis and the neural network autoencoder model.
 4.1 Principal component analysis
-In the SVD analysis, no assumption was used on the statistical properties for the entries of the matrix $\mathbf{X}$.
-In practical applications, the row $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, of this matrix is a measurement sample of a phenomenon.
-Let us assume that the $n$-dimensional vector $\mathbf{x}$ is an observation of a random vector $\mathcal{X}$ which entries are the variables of the phenemenon under study.
+In the SVD, no assumption is used on the statistical properties for the entries of the matrix $\mathbf{X}$.
+In practical applications, the entries of the row $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, are the measurements associated with $n$ attributes or features.
+Let us assume that the $n$-dimensional vector $\mathbf{x}$ is a sample of a random vector $\mathcal{X}$.
 The mean and the covariance of this random vector are: $\mathbb{E}[\mathcal{X}]\triangleq\boldsymbol{\mu}$ and $\mathbb{V}[\mathcal{X}]=\mathbb{E}[(\mathcal{X}-\boldsymbol{\mu})(\mathcal{X}-\boldsymbol{\mu})^{T}]\triangleq\boldsymbol{\Sigma}$.
-The diagonal entries of the matrix $\boldsymbol{\Sigma}$ are the variances of the $\mathcal{X}$ entries: $cov(\mathbf{x}_{j},\mathbf{x}_{j})=\sigma_{j}^{2}$, $j=1,2,\ldots,n$.
-The principal component analysis (PCA) consists of the eigendecomposition of the covariance matrix, followed by a representation of the observations and variables in the reduced space axes, [19, 20].
-Very often, the random vector $\mathcal{X}$ is assumed to be of zero mean.
-When the entries of $\mathcal{X}$ are heterogenous, there can be a large gap between the amplitudes of the variances $\sigma_{j}^{2}$, $j=1,2,\ldots,n$.
-In this case the role of the variables with small variances is hidden in the interpretation of the PCA results.
-To have homogeneous variance data, the centered value of each entry should be divided by the corresponding standard deviation $\sigma_{j}$.
-The covariance matrix becomes a correlation matrix which has only ones as diagonal entries.
-3.1.1 PCA for finite number of observations
-When $m$ observations or objects are available, only estimates can be obtained for the mean vector $\boldsymbol{\mu}$ and the covariance matrix $\boldsymbol{\Sigma}$.
-The sample mean $\bar{\mathbf{x}}$ and the sample variance $s_{j}^{2}$ of the $j$-th entry of the observations $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, are estimated as:
+The diagonal entries of the matrix $\boldsymbol{\Sigma}$ are the attributes variances: $cov(\mathbf{x}_{j},\mathbf{x}_{j})=\sigma_{j}^{2}$, $j=1,2,\ldots,n$.
+The principal component analysis (PCA) consists of the EVD of the covariance matrix, [30, 31].
+When the $\mathcal{X}$’s entries are heterogenous, there can be a large gap between the amplitudes of the variances $\sigma_{j}^{2}$, $j=1,2,\ldots,n$, leading a misinterpretation of the PCA result.
+To have homogeneous variances, the samples are standardized and the correlation matrix is used in the PCA.
+PCA for finite number of samples
+When $m$ samples are available, the sample mean $\bar{\mathbf{x}}$ and the sample variance $s_{j}^{2}$ of the $j$-th entry of the sample $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, are estimated:
 $\bar{\mathbf{x}}=\frac{1}{m}\sum_{i=1}^{m}\mathbf{x}_{i}\mbox{ ; }s_{j}^{2}=\frac{1}{m-1}\sum_{i=1}^{m}(x_{ij}-\bar{x}_{j})^{2}\mbox{ , }j=1,2,\ldots,n$
-From the entries $x_{ij}$ of the matrix $\mathbf{X}$, the entries $x_{ij}^{c}$ of the centered matrix $\mathbf{X}_{c}$ and the entries $x_{ij}^{s}$ of the standardized matrix $\mathbf{X}_{s}$ are obtained as ($i=1,2,\ldots,m$, $j=1,2,\ldots,n$):
-$\mathbf{X}_{c}$ ::: $x^{c}_{ij}=x_{ij}-\bar{x}_{j}$
-$\mathbf{X}_{s}$ ::: $x^{s}_{ij}=(x_{ij}-\bar{x}_{j})/s_{j}$
-Estimates for the sample covariance and the sample variance matrices are obtained using $\mathbf{X}_{c}^{\mathsf{T}}\mathbf{X}_{c}/m$ and $\mathbf{X}_{s}^{\mathsf{T}}\mathbf{X}_{s}/m$, respectively.
-Using the expressions between the SVD factors and those of the eigendecomposition, the PCA result can be obtained from the SVD results of the matrix $\mathbf{X}_{c}$ or the matrix $\mathbf{X}_{s}$.
-3.1.2 Representation of objects or variables in reduced space
-The matrices $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and $\mathbf{XX}^{\mathsf{T}}$ have the same eigenvalues and related eigenvectors. The eigenvalue $\lambda$ allows to write:
+Then, the entries $x_{ij}^{c}$ of the centered matrix $\mathbf{X}_{c}$ and the entries $x_{ij}^{s}$ of the standardized matrix $\mathbf{X}_{s}$ are obtained as ($i=1,2,\ldots,m$, $j=1,2,\ldots,n$):
+$\mathbf{X}_{c}$ $:$ $x^{c}_{ij}=x_{ij}-\bar{x}_{j}$
+$\mathbf{X}_{s}$ $:$ $x^{s}_{ij}=(x_{ij}-\bar{x}_{j})/s_{j}$
+Estimates for the sample covariance and the sample variance matrices are obtained as $\mathbf{X}_{c}^{\mathsf{T}}\mathbf{X}_{c}/m$ and $\mathbf{X}_{s}^{\mathsf{T}}\mathbf{X}_{s}/m$, respectively.
+Using the expressions between the SVD factors and those of the EVD, the PCA result can be obtained by using the matrix $\mathbf{X}_{c}$ or the matrix $\mathbf{X}_{s}$.
+Representation of the samples or the attributes in a reduced space
+The matrices $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and $\mathbf{XX}^{\mathsf{T}}$ have the same eigenvalues and related eigenvectors:
 $\mathbf{X}^{\mathsf{T}}\mathbf{Xv}$ $=$ $\lambda\mathbf{v}$ (26)
@@
 Let $r$ be the reduced dimension used in the PCA, typically $r=2$ or $3$ for a visual representation.
-From the PCA results of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$, expressions for the objects in the $r$-dimension variable space are $\mathbf{X}\mathbf{V}_{r}$, where $\mathbf{V}_{r}$ is a matrix with the $r$ largest eigenvectors.
-Using the expression for the matrix $\mathbf{U}$ in relation (14), expressions for the variables in the $m$-dimensional space are obtained as: $\mathbf{X}^{\mathsf{T}}\mathbf{XV}_{r}\boldsymbol{\Lambda}_{r}^{-1/2}$, where $\boldsymbol{\Lambda}_{r}^{-1/2}$ is the inverse of the diagonal matrix obtained using the $r$ largest nonzero eigenvalues.
-The PCA results plot for the iris data is shown in Figure 2.
-Figure 2: Plot of the iris observations in the first and third principal components axes.
+From the PCA result of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$, expressions for the samples in the $r$-dimension attribute space are $\mathbf{X}\mathbf{V}_{r}$, where $\mathbf{V}_{r}$ is a matrix with the $r$ first largest eigenvectors.
+Using the expression for the matrix $\mathbf{U}$ in the relation (6), the attributes are expressed in the $m$-dimensional space as: $\mathbf{X}^{\mathsf{T}}\mathbf{XV}_{r}\boldsymbol{\Lambda}_{r}^{-1/2}$, where $\boldsymbol{\Lambda}_{r}^{-1/2}$ is the square root inverse of the diagonal matrix associated with the $r$ first largest nonzero eigenvalues.
+The PCA result plot for the iris data is shown in Figure 2 where the $r=3$ first singular values and singular vectors were computed, the default settings were used for the other parameters.
+Figure 2: Plot of the iris samples in the first and third principal components axes.
 4.2 Autoencoder
-An autoencoder (AE) is a neural network (NN) model which allows to encode input, store the result internally, and finally decode the stored result to obtain an output.
+An autoencoder is a neural network (NN) model which allows to encode input, store the result internally, and finally decode the stored result to obtain an output.
 The output is not necessary strictly equal to the input.
 In comparison, the SVD leads to the matrices $\mathbf{U}$, $\mathbf{D}$ and $\mathbf{V}$ which are an internal representation of the input matrix $\mathbf{X}$.
-Then, using a rank $r$ value, an approximation of the input is obtained: $\hat{\mathbf{X}}=\mathbf{U}_{r}\mathbf{D}_{r}\mathbf{W}_{r}^{\mathsf{T}}$.
-This approximation matrix is of rank $r$ and the SVD uses linear operations to solve the dimensionality reduction problem or data compression.
+Then, using the $r$ first largest singular values, a low-rank approximation is obtained: $\hat{\mathbf{X}}=\mathbf{U}_{r}\mathbf{D}_{r}\mathbf{W}_{r}^{\mathsf{T}}$.
+Thus, the SVD uses linear algebra operations to solve the dimensionality reduction problem or data compression.
 The data can also be compressed using a feedforward NN model with few layers, [49, 8, 35].
-A NN layer consists of a weighted combination of an input, $\mathbf{x}$, followed by a transformation function $\varphi(.)$ to obtain an output: $\mathbf{u}=\mathbf{x}^{\mathsf{T}}\mathbf{W}+\mathbf{w}_{o}$; $\mathbf{v}=\varphi(\mathbf{u})$, where $\mathbf{W}$ is weight matrix and $\mathbf{w}_{o}$ is a threshold or bias vector.
-The transformation operates on each entry of the vector $\mathbf{u}$.
+A NN layer is a weighted combination of an input, $\mathbf{x}$, followed by a transformation function $\varphi(.)$ to obtain an output: $\mathbf{u}=\mathbf{x}^{\mathsf{T}}\mathbf{W}+\mathbf{w}_{o}$; $\mathbf{v}=\varphi(\mathbf{u})$, where $\mathbf{W}$ is a weight matrix and $\mathbf{w}_{o}$ is a threshold or bias vector.
+The transformation operates on each entry $u$ of the vector $\mathbf{u}$.
 Three examples of the transformation function are the sigmoid, $\sigma(u)=1/(1+e^{-u})$, the rectified linear unit, $ReLU(u)=max(0,u)$, and the identity $I(u)=u$.
-A two layers feedforward NN model has been used in [22] to perform a nonlinear SVD for obtaining an approximation of the input $\mathbf{X}$.
-Each row of the matrix $\mathbf{X}$ is an $n$-dimensional vector $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, and the output of the encoding layer is obtained as:
+A two layers feedforward NN model has been used in [35] to perform a nonlinear SVD for an approximation of $\mathbf{X}$.
+The encoder output $\mathbf{v}_{i}$ is obtained for each input $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, as:
 $\mathbf{v}_{i}=\sigma\left(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{W}_{e}+\mathbf{w}_{e}\right)$
-where $\mathbf{W}_{e}$ is a weight matrix of size $n\times r$ and $\mathbf{w}_{e}$ is a bias vector of order $n$. The output of the decoding layer is obtained as:
-$\hat{\mathbf{x}}_{i}=\sigma\left(\mathbf{v}_{i}^{\mathsf{T}}\mathbf{W}_{d}+\mathbf{w}_{d}\right)=\sigma\left(\sigma\left(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{W}_{e}+\mathbf{w}_{e}\right)^{\mathsf{T}}\mathbf{W}_{d}+\mathbf{w}_{d}\right)$
-where $\mathbf{W}_{d}$ is a weight matrix of size $r\times n$ and $\mathbf{w}_{d}$ is a bias vector of order $n$.
-The input $\mathbf{x}_{i}$ and its approximation by the decoder output are used to define an $\ell_{2}$ norm:
+where $\mathbf{W}_{e}$ is the encoder weight matrix of size $n\times r$ and $\mathbf{w}_{e}$ is a bias vector of order $r$. The output of the decoding layer is obtained as:
+$\hat{\mathbf{x}}_{i}=\sigma\left(\mathbf{v}_{i}^{\mathsf{T}}\mathbf{W}_{n}+\mathbf{w}_{n}\right)=\sigma\left(\sigma\left(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{W}_{e}+\mathbf{w}_{e}\right)^{\mathsf{T}}\mathbf{W}_{n}+\mathbf{w}_{n}\right)$
+where $\mathbf{W}_{n}$ is a weight matrix of size $r\times n$ and $\mathbf{w}_{n}$ is a bias vector of order $n$.
+The input $\mathbf{x}_{i}$ and its approximation, the decoder output, can be used to define a $\ell_{2}$-norm:
 $\|\mathbf{x}_{i}-\hat{\mathbf{x}}_{i}\|$
-From the $m$ observations, the weight matrices, $\mathbf{W}_{e}$ and $\mathbf{W}_{d}$, and the bias vectors, $\mathbf{w}_{e}$ and $\mathbf{w}_{d}$ are obtained after minimization of the $\ell_{2}$-norm using a gradient search optimization algorithm.
-Instead of using a single layer in the encoder, $k$ layers, $k>1$, can be used to progressively increase the compression rate.
-In this case, an inverse operation is done using $k$ decoding layers.
-The approximation matrix obtained using the feedforward NN model is not necessary of reduced rank even the number of columns of the encoder weight matrix is $r<n$.
-However, equivalent approximation to SVD result can be obtained using the AE model.
-The difference matrix $\mathbf{X}-\hat{\mathbf{X}}$ can be viewed as the noise part of the data.
-Therefore, the approximation $\hat{\mathbf{X}}$ can be used in subsequent supervised or unsupervised data analysis methods.
-3.2.1 Identity transformation function
+From $m$ samples, the weight matrices, $\mathbf{W}_{e}$ and $\mathbf{W}_{n}$, and the bias vectors, $\mathbf{w}_{e}$ and $\mathbf{w}_{n}$ are obtained after minimization of the $\ell_{2}$-norm using a gradient search optimization algorithm.
+The approximation matrix obtained using a nonlinear activation feedforward NN model is not necessary of reduced rank even the number of columns of the encoder weight matrix is $r<n$.
+However, equivalent approximation to SVD result can be obtained using the autoencoder model.
+The difference matrix $\mathbf{X}-\hat{\mathbf{X}}$ can be viewed as the noise part of the data and the approximation $\hat{\mathbf{X}}$ can be used in subsequent supervised or unsupervised data analysis methods.
+Identity transformation function
 The identity transformation function has been used in [52, 55].
-The transpose of the matrix $\mathbf{X}$ is used in [33] leading to $\|\mathbf{X}-\mathbf{WW}^{\mathsf{T}}\mathbf{X}\|$ while the constrained optimization objective function of the relation (16) is used in [36, Chapter 7].
-In these two references, the classical gradient search optimization method is used to obtain the matrix $\mathbf{W}$.
-The algorithm 1 solves the same problem by converting the gradient search into a power method algorithm and adds a convergence acceleration through the choice of the parameter $q$.
-When an estimation is available for the weight matrix $\mathbf{W}$, then an proximation of the input matrix $\mathbf{X}$ is: $\hat{\mathbf{X}}=\mathbf{XWW}^{\mathsf{T}}$.
-Therefore, an approximation of the input vector $\mathbf{x}_{i}$ is: $\hat{\mathbf{x}}_{i}=\mathbf{x}_{i}^{\mathsf{T}}\mathbf{WW}^{\mathsf{T}}$.
-MNIST data.
-For illustration, the MNIST data, [26], was used.
-This dataset consists of $60,000$ training images and $10,000$ test images.
-Each image of size $28\times 28$ is an observation or vector of dimension $784$.
+The transpose of the matrix $\mathbf{X}$ is used in [52] leading to $\|\mathbf{X}-\mathbf{WW}^{\mathsf{T}}\mathbf{X}\|$ while the constrained optimization objective function of the relation (16) is used in [55, Chapter 7].
+In these references, the classical gradient search optimization method is used to obtain the matrix $\mathbf{W}$.
+The $k$ columns of this matrix are not necessary the $k$ first singular vectors of the data matrix.
+For the proposed method, the $k$ first largest singular vectors are the $k$ columns of the autoencoder model weight matrix.
+When an estimation is available for $\mathbf{W}$, then approximations for the matrix $\mathbf{X}$ and the vector $\mathbf{x}_{i}$ are $\hat{\mathbf{X}}=\mathbf{X}\hat{\mathbf{W}}\hat{\mathbf{W}}^{\mathsf{T}}$ and $\hat{\mathbf{x}}_{i}=\mathbf{x}_{i}^{\mathsf{T}}\hat{\mathbf{W}}\hat{\mathbf{W}}^{\mathsf{T}}$, respectively.
+MNIST data
+This dataset consists of $60,000$ training images and $10,000$ test images, [40].
+Each image of size $28\times 28$ is a sample, a vector of order $784$.
 The entries of this vector are pixels which values are from the set $\{0,1,\ldots,255\}$.
@@
 The last $72$ singular values are zero.
-The total information captured by the first $50$, $100$ or $256$ largest singular values are $45.41\%$, $60.45\%$ and $82.63\%$, respectively.
+The total information captured by the $50$, $100$ or $256$ first largest singular values are $45.41\%$, $60.45\%$ and $82.63\%$, respectively.
 (a) log scale singular values
 (b) cumulative reconstruction rate
-Figure 3: Log scale singular values and cumulative percentage reconstruction rate
-The largest $256$ singular values were computed using the training images data.
-The weight matrix obtained was used to estimate approximates for the first $5$ test images.
-The results are shown in Figure 4, where the first column is the test image and the $3$ following columns are for the approximation using the first $50$, $100$ or $256$ singular values, respectively.
+Figure 3: Log scale singular values and cumulative percentage reconstruction rate for the MNIST images
+The $256$ first largest singular values were computed using the training images data and each method.
+The results are shown in Table 5.
+In the first part of this Table, the default parameter itmax (200) is used while itmax = 150(a), 100(b) and 50(c), are used for the 3 last rows results (proposed method only).
+Like for the trachea dataset, the rsvd method has the highest MSE value, with a reconstruction rate almost similar to that of the other methods.
+When accurate singular values are not required, a fast runtime is obtained by using a small value for the parameter itmax in the algorithm 1.
+Table 5: Calculation of the $256$ first singular values for the MNIST images data using 6 methods: runtime, reconstruction rate and MSE comparison results
+method runtime reconst.rate MSE
+svd 85.26 82.63 -
+propack.svd 71.35 82.63 <<1e-10
+RSpectra 47.50 82.63 <<1e-10
+irlba 219.74 82.63 5.99e-6
+rsvd 87.68 82.35 16,631.1
+psvd 82.75 82.63 29.16
+psvd(a) 69.72 82.63 71.43
+psvd(b) 57.71 82.63 173.8
+psvd(c) 45.35 82.63 2,777.4
+The weight matrix obtained using the proposed method was used to have the $5$ first test images approximations.
+The results are shown in Figure 4, where the first column is the test image and the $3$ following columns are the results using the $50$, $100$ or $256$ first largest singular values approximations, respectively.
 The compression rates for these approximations are $15.68$, $7.84$ and $3.06$, respectively.
-Figure 4: First 5 test images, from left to right: input, use of the $50$, $100$ or $256$ largest singular values to obtain approximation.
+Figure 4: 55 first test images, from left to right: input, use of the $50$, $100$ or $256$ first largest singular values to obtain approximation.
 5 Conclusion
-A new method is proposed for computing the singular values of a matrix.
-This method allows to compute all or a subset of the largest singular values and associated vectors.
-It can be applied to small or high dimensional dataset.
-The accuracy of the results depends on the user tolerance setting.
+A new method is proposed for computing the singular values of a real-valued entries matrix.
+This method consists of using a gradient search optimizer to minimize a Frobenius norm and allows to compute all or a subset of the $k$ first largest singular values and singular vectors.
+The accuracy of the results depends on the user settings.
 More accuracy may necessitate a long runtime depending on the behavior of the singular values decay.
-The proposed method especially consists of searching for an eigendecomposition of a square symmetric matrix by using a gradient method to minimize a Frobenius norm.
+A parallel implementation may allow to speed up the runtime of the proposed SVD calculation algorithm.
 The method was applied to principal component analysis and to neural network autoencoder model.
+An independent optimization method is often used in the literature to get the singular vectors which are not necessary the $k$ first columns of the encoder weight matrix unlike for the method proposed.
 Annexes
 The proposed method is implemented in the psvd R package available on comprehensive R archive network [54]: http://cran.r-project.org/web/packages/psvd/.
-In the first version, the power parameter $q$ is fixed to $2$ and only a sample from standard normal distribution is used as initial projection matrix.
+The power parameter $q$ can be chosen in the set $\{2,3,4\}$, a sample from standard normal distribution is used as initial projection matrix.
 The basic power method algorithm is given below to show the similarity with the proposed algorithm.
@@
 A C code is used for this part in the psvd R package.
-Reader can use her/his preferred computing language to implement all these algorithms.
+Reader can use her/his preferred computing environment to implement these algorithms.
 Power method
 Data: square matrix $\mathbf{A}$
-Result: orthonormal vector $\mathbf{v}$
+Result: orthogonal vector $\mathbf{v}$
 1 Initialization: choose an initial vector $\mathbf{v}^{(0)}$ and compute $\mathbf{v}^{(0)}\leftarrow\mathbf{v}^{(0)}/\|\mathbf{v}^{(0)}\|$, set tolerance $\epsilon$;